### v1

# Training Large Neural Networks With Low-Dimensional Error Feedback

Abstract
Training deep neural networks typically relies on backpropagating high-dimensional error signals—a computationally intensive process with little evidence supporting its implementation in the brain. However, since most tasks involve low-dimensional outputs, we propose that low-dimensional error signals may suffice for effective learning. To test this hypothesis, we introduce a novel local learning rule based on Feedback Alignment that leverages indirect, low-dimensional error feedback to train large networks. Our method decouples the backward pass from the forward pass, enabling precise control over error signal dimensionality while maintaining high-dimensional representations. We begin with a detailed theoretical derivation for linear networks, which forms the foundation of our learning framework, and extend our approach to nonlinear, convolutional, and transformer architectures. Remarkably, we demonstrate that even minimal error dimensionality—on the order of the task dimensionality—can achieve performance matching that of traditional backpropagation. Furthermore, our rule enables efficient training of convolutional networks, which have previously been resistant to Feedback Alignment methods, with minimal error. This breakthrough not only paves the way toward more biologically accurate models of learning but also challenges the conventional reliance on high-dimensional gradient signals in neural network training. Our findings suggest that low-dimensional error signals can be as effective as high-dimensional ones, prompting a reevaluation of gradient-based learning in high-dimensional systems. Ultimately, our work offers a fresh perspective on neural network optimization and contributes to understanding learning mechanisms in both artificial and biological systems.
1.   Introduction
The brain, despite its vast complexity, operates within the constraints of a low-dimensional world. Every movement—lifting a hand, turning a head—follows finite joint constraints. Navigation unfolds in modest spatial dimensions, and object recognition is drawn from a limited set of familiar categories. Even when possibilities seem vast, they are dwarfed by the sheer number of neurons and synapses in any brain region. Artificial neural networks reflect this paradox: the number of neurons in each layer often far exceeds the task-relevant variables. This contrast between task and internal representation dimensionality challenges us to rethink how we train such large networks and explore new learning paradigms.
Modern deep learning methods rely on high-dimensional gradients that propagate backward through layers, ensuring a precise credit assignment to each parameter [1, 2]. However, this process is computationally intensive and biologically implausible, given the lack of direct evidence for backpropagation in the brain [3, 4]. If the loss gradient does not propagate back through the network, it must be obtained through alternative routes [5]. In this case, there is no reason to assume that each neuron receives detailed information about the local gradient. Moreover, the brain likely operates with low-dimensional error signals, reflecting the inherently low-dimensional nature of tasks. This raises an essential question: Can large neural networks be trained efficiently with constrained, low-dimensional error signals? If error dimensionality is tied to task complexity rather than network size, it could revolutionize how we train large networks and improve our understanding of biological learning.
Despite the importance of this question, the effect of low-dimensional error feedback on learning and neural representations was never properly explored. The strong coupling of feedforward and backward passes in backpropagation, where both processes use the same synapses, has limited investigations into alternative training methods. While approaches such as Feedback Alignment have been proposed [5], the necessity of full gradients for effective learning has not been questioned.
In this study, we propose that high-dimensional error propagation is not strictly necessary for training deep, overparameterized networks. We introduce a novel learning rule based on Feedback Alignment that decouples forward and backward passes, enabling precise control over the dimensionality of error signals. Our method exploits the low-dimensional nature of real-world tasks to train large networks efficiently without compromising their representational capacity, achieving performance on par with backpropagation. Our theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance. Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is both feasible and universal. Finally, we explore how error dimensionality influences neural representations by examining its effects on receptive fields in an early visual system model, underscoring the broader implications of our findings.
2.   Background and related work
We consider a multilayered perceptron with $L$ layers, each layer $l$ computes its output as $\boldsymbol{h}_{l}=f(W_{l}\boldsymbol{h}_{l-1})$, where $W_{l}$ is the weight matrix and $f$ is an element-wise activation function. The input to the network is $\boldsymbol{h}_{0}=\boldsymbol{x}$, and the final network output is $\hat{\boldsymbol{y}}=f_{L}(W_{L}\boldsymbol{h}_{L-1})$, which approximates the target $\boldsymbol{y}$. The task dimensionality, denoted $d$, is at most the number of components in $\boldsymbol{y}$ and $\hat{\boldsymbol{y}}$. Training the network involves minimizing a loss function $\mathcal{L}(\boldsymbol{y},\hat{\boldsymbol{y}})$ by adjusting the weights $\{W_{l}\}$. The error signal at the output layer, $\delta_{L}=\frac{\partial\mathcal{L}}{\partial\hat{\boldsymbol{y}}}$, is a $d$-dimensional vector, typically much smaller than the number of neurons in the hidden layers.
Backpropagation (BP) [1] is the standard approach for training neural networks. It propagates the error backward through the network using $\delta_{l}=W_{l+1}^{T}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1})$, and updates the weights using $\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{T}$, where $\eta$ is the learning rate. However, this method requires an exact transposition of the forward weights, $W_{l+1}^{T}$, which is biologically implausible [6, 7]. Moreover, BP tightly couples the error propagation with the forward pass, limiting the ability to explore how different properties of the error signal affect learning dynamics.
Feedback Alignment (FA) [5] was proposed to address the biological limitations of BP by replacing $W_{l+1}^{T}$ with a fixed random matrix $B_{l}$. The error is computed as:
$\delta_{l}=B_{l}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1}),$ (1)
decoupling the forward and backward weights and providing a more biologically plausible mechanism. However, FA struggles to scale effectively in deep architectures, such as convolutional neural networks (CNNs), where it often fails to achieve competitive performance [8].
An extension of FA involves adapting $B_{l}$ by updating it alongside the forward weights $W_{l}$ to improve their alignment [9, 10]:
$\Delta B_{l}=-\eta\boldsymbol{h}_{l-1}\delta_{l}^{T}-\lambda B_{l},\quad\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{T}-\lambda W_{l},$ (2)
where $\lambda$ is a regularization parameter. Although this adaptive approach improves performance by better aligning forward and backward weights, it still requires high-dimensional error signals and struggles to match BP performance in complex architectures like CNNs. Furthermore, in Section 3, we show that this approach fails when the matrix $B$ is low-rank and the dimensionality of the error is constrained.
Other studies have explored the use of fixed sparse feedback matrices to reduce the dimensionality of error propagation [11]. However, these approaches result in significantly lower performance and do not provide a systematic framework for studying how error constraints affect learning and representation formation.
Beyond FA-based methods, several studies have shown that weight updates using backpropagation can result in a low-dimensional weight update [12, 13, 14] and favor low-rank solutions [15]. These findings support our hypothesis that low-dimensional feedback is sufficient to train deep networks. However, no previous work has considered training with a constrained error pathway, and the effects of error dimensionality and training have not been systematically studied.
In this work, our aim is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of neural networks. To this end, we introduce a novel learning scheme, Restricted Adaptive Feedback (RAF), that allows flexible control over the dimensionality of the errors (Figure 1).
Our main contributions are:
  1. 1.
    We present a novel learning rule, Restricted Adaptive Feedback (RAF), which matches BP performance while requiring minimal error signals. We provide a detailed derivation of the learning dynamics in a simple linear case, establishing a foundational understanding of how RAF operates.
  2. 2.
    We demonstrate that nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals, highlighting the versatility of RAF in practical scenarios.
  3. 3.
    We show that more complex yet highly useful architectures, such as convolutional networks and transformers, can also be effectively trained with low-dimensional feedback.
  4. 4.
    We reveal that error dimensionality shapes the receptive fields in a model of the ventral visual system, offering new insights into the relationship between learning mechanisms and biological neural representations.
In the final section, we discuss the broader implications of our results for both neuroscience and machine learning.
Figure 1: Illustration of different approaches for propagating error to hidden layers. From left to right: Backpropagation (BP) propagates error using the exact transpose of the forward weights. Feedback Alignment (FA) replaces the transposed weights with fixed random feedback matrices, which gradually align with the forward weights during training. Restricted Adaptive Feedback (RAF) constrains error propagation through low-rank feedback matrices, limiting error dimensionality and preventing an exact mirroring of BP. Direct Restricted Adaptive Feedback (dRAF) extends RAF by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
3.   Low-dimensional error feedback in linear networks
We begin our analysis by studying learning dynamics in multilayered linear networks. Although linear models may seem overly simplistic, they can exhibit rich learning dynamics due to the nonlinearity introduced by the loss function [16]. Additionally, imposing dimensional constraints on linear networks yields insightful results that extend beyond the linear case.
A linear problem
We consider a simple linear transformation problem with a low-dimensional structure, $\boldsymbol{y}=A\boldsymbol{x}$. Here, $\boldsymbol{x}\in\mathbb{R}^{n}$ represents the $n$-dimensional input, and $\boldsymbol{y}\in\mathbb{R}^{m}$ represents the target. The matrix $A$ is a rank-$d$ matrix defined as $A=\sum_{j=1}^{d}\boldsymbol{u}_{j}\boldsymbol{v}_{j}^{T}$, where $\boldsymbol{u}_{j}\in\mathbb{R}^{n}$ and $\boldsymbol{v}_{j}\in\mathbb{R}^{m}$ are random Gaussian vectors, and we assume $d\ll n$. Our data set consists of $p$ training samples $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}_{\mu=1}^{p}$, with each input vector $\boldsymbol{x}^{\mu}$ being i.i.d. according to the standard normal distribution $\boldsymbol{x}^{\mu}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{1})$. The labels are given by $\boldsymbol{y}^{\mu}=A\boldsymbol{x}^{\mu}+\boldsymbol{\xi}^{\mu}$, where $\boldsymbol{\xi}^{\mu}$ is additive Gaussian noise with zero mean and unit variance.
The goal is to learn the low-dimensional structure of $A$ from the $p$ samples using a linear neural network. For simplicity, we assume that $p$ is sufficiently large to allow the network to fully recover the structure of $A$.
A linear network model.
To study the effects of restricted error pathways, we consider a simple linear network with three layers: an input layer $\boldsymbol{x}\in\mathbb{R}^{n}$, a hidden layer $\boldsymbol{h}\in\mathbb{R}^{k}$, and an output layer $\boldsymbol{y}\in\mathbb{R}^{m}$. The input and hidden layers are connected by the weight matrix $W_{1}\in\mathbb{R}^{k\times n}$, and the hidden and output layers are connected by the weight matrix $W_{2}\in\mathbb{R}^{m\times k}$. The output of the network can be expressed as $\boldsymbol{y}=W_{2}W_{1}\boldsymbol{x}$ (Figure 2).
The network is trained to minimize the quadratic empirical loss function:
$L=\frac{1}{p}\sum_{\mu}\|\boldsymbol{y}(\boldsymbol{x^{\mu}})-W_{2}W_{1}\boldsymbol{x}^{\mu}\|^{2}.$ (3)
We apply Feedback Alignment (FA) to update $W_{1}$, which does not have direct access to the loss gradient. Instead of backpropagating the error through $W_{2}^{T}$, we use a fixed low-rank feedback matrix $B$. This provides an alternative pathway for propagating the error signal to $W_{1}$.
For a given data point $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}$, the weight updates, derived from the FA framework, are given by:
$\Delta W_{1}^{\mu}$ $=\eta B^{T}(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
$\Delta W_{2}^{\mu}$ $=\eta(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T}W_{1}^{T}.$
Here, $\eta$ represents the learning rate, and the update for $W_{1}$ is computed using the indirect error feedback provided by $B$, while $W_{2}$ receives the full error signal directly from the output.
Constraining error dimensionality with low-rank feedback
To control the dimensionality of the error feedback, we impose a low-rank constraint on the feedback matrix $B$. Rather than allowing full-dimensional feedback, we decompose $B$ as $B=QP$, where $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$. When $r<\min(k,m)$, $B$ is low rank, which means that it can project the error signal onto at most $r$ independent directions.
This low-rank structure introduces an "$r$-bottleneck," which limits the flow of error information. In linear settings, the problem is solvable using a single weight matrix, rendering the training of $W_{1}$ unnecessary. However, backpropagation dynamics still adjust these weights  [16]. By controlling the value of $r$, we can systematically study how reducing the dimensionality of the error signal impacts learning. Initially, we follow the original Feedback Alignment framework, keeping $Q$ and $P$ as random matrices. However, as we will demonstrate, allowing $Q$ and $P$ to learn is crucial to high performance.
3.1 Learning dynamics
Our analysis extends the framework established by Saxe et. al. [16] to incorporate indirect feedback with constrained dimensionality. We begin by characterizing the task across the $p$ data points through the input-output covariance matrix, $\Sigma_{io}=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}(\boldsymbol{x}^{\mu})^{T}$, which captures the correlation between input vectors $\boldsymbol{x}$ and output vectors $\boldsymbol{y}$. Performing Singular Value Decomposition (SVD), we obtain $\Sigma_{io}=USV^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ contain the left and right singular vectors, respectively, and $S\in\mathbb{R}^{m\times n}$ is a rectangular diagonal matrix of singular values. For sufficiently large $p$, the first $d$ singular values in $S$ correspond to the prominent directions in the data (i.e., the singular values of $A$), while the remaining singular values are $O(1/\sqrt{p})$ and dominated by noise.
To track how training aligns the network weights with these prominent directions, we rotate the weight matrices $W_{1}$, $W_{2}$, and $B$ using the singular vectors of $\Sigma_{io}$. This transformation simplifies the analysis by aligning the network’s weight dynamics with the key data directions:
$W_{1}=\bar{W}_{1}V^{T},\quad W_{2}=U\bar{W}_{2},\quad B=\bar{B}U^{T},$
where $\bar{W}_{1}$, $\bar{W}_{2}$, and $\bar{B}$ represent the transformed weight matrices. This rotation aligns the weight dynamics with the dominant singular vectors, allowing us to focus on how the network captures the important features of the data.
Since the inputs are uncorrelated, we can apply these transformations to the iterative weight-update equations derived from the FA learning rule. Assuming a small learning rate $\eta\ll 1$ with full-batch updates, we express the weight updates in continuous time:
$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}^{T}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
where $\eta=dt/\tau$. This continuous form captures the dynamics of the learning process, allowing us to study it from a dynamical systems perspective. By analyzing these equations, we can identify fixed points and evaluate their stability, providing insight into how the network converges and learns under constrained feedback. Figure 2, shows how the singular vectors of $W_{2}W_{1}$ align to the corresponding singular vectors of $\Sigma_{io}$
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$, with $r=64$ and $r=8$, respectively. The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$. (d, e) Same as (b, c) but for Restricted Adaptive Feedback (RAF), where both $Q$ and $P$ are trained. (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully. (g) Training $P$ to align with the error, as in RAF, ensures full recovery of the top $r$ singular components, improving convergence and alignment.
Stationary solutions for the training.
Training halts when the right-hand side of the weight update equations (5) vanishes, indicating that the dynamics have reached a stable fixed point. At this fixed point, the update equation for $\bar{W}_{1}$ leads to the condition:
$\bar{B}(S-\bar{W}_{2}\bar{W}_{1})=0\implies S_{jj}\boldsymbol{B}_{:,j}=\sum_{i=1}^{m}\boldsymbol{B}_{:,i}(\bar{W}_{2}\bar{W}_{1})_{ij}\quad\forall j,$ (6)
where $\boldsymbol{B}_{:,j}\in\mathbb{R}^{k}$ is the $j$-th column of $\bar{B}$, and $S_{jj}$ is the $j$-th singular value of $\Sigma_{io}$. This equation indicates that, at the stationary point, the weight products $\bar{W}_{2}\bar{W}_{1}$ must align with the singular modes of the data.
However, since $B$ is of rank $r$, the feedback matrix $\bar{B}$ can only span at most $r$ independent directions. If $r=m$, the system has enough feedback dimensionality to align perfectly with the singular values in $S$, recovering the full structure of $\Sigma_{io}$ as demonstrated in Saxe et. al. [16]. In this case, the training successfully converges to a unique solution where $\bar{W}_{2}\bar{W}_{1}=S$ (Figure 2b).
Crucially, when $r<m$, the feedback matrix $\bar{B}$ lacks the sufficient rank to fully capture the $m$ independent directions in $S$. As a result, eq. (6) becomes under-determined, leading to potentially infinite solutions. This means that the trained weights $\bar{W}_{2}\bar{W}_{1}$ may not align with the true data structure encoded in $\Sigma_{io}$ (Figure 2c)
The implications of the uderdetrimed fixed point solutions are surprising. Naively, one would expect that if the bottleneck is not too narrow, the projections of the error would maintain the necessary structure to guide the learning. Specifically, the Johnson–Lindenstrauss lemma  [17] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal would be maintained. Nevertheless, our analysis shows that the solutions weight are not guaranteed to converge to the correct solution. Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail. The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.
3.2 Training the feedback weights
From the previous analysis, it is apparent that the feedback matrix $B$ must be trained for the learning dynamics to converge to the correct solution of Eq. (6). When $B$ is low-rank and fixed, the network lacks sufficient capacity to transmit the full error information, which can impede learning.
One approach is to adopt a learning rule inspired by Kolen-Pollack, as in eq. (2). When $B=QP$, the updates are given by
$\Delta W_{2}^{\mu}=\eta\boldsymbol{\delta}^{\mu}\boldsymbol{x}^{\mu T}W_{1}^{T}-\lambda W_{2}\quad\text{ and }\quad\Delta Q^{\mu}=\eta W_{1}\boldsymbol{x}^{\mu}\boldsymbol{\delta}^{\mu T}P^{T}-\lambda Q,$ (7)
where $\boldsymbol{\delta}^{\mu}=\frac{\partial\mathcal{L}(\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu})}{\partial\boldsymbol{y}}=\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}$ is the error gradient for the $\mu$-th data point, $\eta$ is the learning rate, and $\lambda$ is the regularization parameter. Importantly, in this framework only the column space of $B$, or in our case $Q$ are updated.
However, adopting this learning framework is insufficient when $P$ remains fixed and randomly initialized (Figure 2c). Since $P$ can project onto at most $r$ unique directions of the error, it may not align with the relevant error subspace. In this case, the network may converge to an incorrect solution, as indicated by the non-uniqueness of solutions to (6) when $m>r$.
Ideally, we want $P$ to be an orthogonal matrix whose $r$ columns span the top $r$ principal directions of the output-output correlation matrix $\Sigma^{yy}$. This alignment ensures that the most significant components of the error are propagated back through the network.
In cases where the output correlations are unknown, we can update $P$ using a modified Oja learning rule [18]:
$\Delta P^{\mu}=\eta P\boldsymbol{y}^{\mu}\boldsymbol{y}^{\mu T}(I-P^{T}P)-\lambda P,$ (8)
where $I$ is the identity matrix. This rule adjusts $P$ incrementally so that its columns converge to the top $r$ principal components of the outputs $\{\boldsymbol{y}^{\mu}\}$.
By training both $Q$ and $P$, we allow the feedback matrix $B=QP$ to adaptively align with the relevant error directions, enabling the network to learn the correct mappings even under constrained feedback dimensionality.
Repeating the linear analysis that led to (5), we extend the derivation to our case with adaptive feedback weights. We define the transformed feedback matrix as $P=\bar{P}U^{T}$. This transformation aligns the feedback matrix $P$ with the principal components of the data, simplifying the analysis.
Taking the continuous-time limit (with $\eta\to 0$ and $\eta p=\tau$), we obtain a set of differential equations that describe the learning dynamics of the forward and backward weights:
$\tau\frac{d\bar{W}_{1}}{dt}$ $=\bar{B}^{T}(S-\bar{W}_{2}\bar{W}_{1}),$ (9)
$\tau\frac{d\bar{W}_{2}}{dt}$ $=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T}-\lambda\bar{W}_{2},$
and
$\tau\frac{dQ}{dt}$ $=\bar{W}_{1}(S-\bar{W}_{2}\bar{W}_{1})^{T}\bar{P}^{T}-\lambda Q,$ (10)
$\tau\frac{d\bar{P}}{dt}$ $=\bar{P}SS^{T}(I-\bar{P}^{T}\bar{P})-\lambda\bar{P}.$
Here, $\bar{B}=Q\bar{P}$ and $\lambda$ is the regularization parameter. Note that $Q$ is not affected by rotation, as it does not come in contact with either the input or the output. The full derivation can be found in the Appendix.
Notably, the updates to the feedback weights are local and follow learning rules that were well-studied in theoretical neuroscience [18, 19, 20, 21]. Local plasticity makes our framework an attractive alternative for backpropagation in models of brain circuits.
We refer to this learning framework as Restricted Adaptive Feedback (RAF). Figure 2 compares the learning dynamics of RAF with those of BP and FA where only the $Q$ matrix is learned. The results demonstrate that RAF effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality. Furthermore, RAF will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
3.3 Restricted adaptive feedback in deep architectures
Our weight-update equations can be naturally extended to deeper networks. In the single hidden layer model above, the backward weights $P$ were updated using the true labels $\boldsymbol{y}^{\mu}$. However, in deeper models, hidden layers do not have ground-truth representations. Instead, each layer relies on the local error signal, which propagates through the network according to $\boldsymbol{\delta}_{l}^{\mu}=Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}$. This error signal $\boldsymbol{\delta}_{l}^{\mu}$ provides the necessary information for learning at the layer $l$.
To update the feedback weights $P_{l}$ in the absence of ground-truth representations, we use local error signals $\boldsymbol{\delta}_{l+1}^{\mu}$. As in the single-layer model, we use Oja’s rule to adjust $P_{l}$ to span the principal components of the error at the next layer, $\boldsymbol{\delta}_{l+1}^{\mu}$, ensuring efficient error propagation.
The complete update rules for layer $l$ are given by
$\Delta W_{l}^{\mu}&=\eta Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{h}_{l}^{\mu T}-\lambda W_{l}\end{aligned}\qquad\text{and}\qquad\begin{aligned} \Delta Q_{l}^{\mu}&=\eta\boldsymbol{h}_{l}^{\mu}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}-\lambda Q_{l},\\ \Delta P_{l}^{\mu}&=\eta P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu T}(I-P_{l}^{T}P_{l})-\lambda P,\end{aligned}$ (11)
where $\eta$ is the learning rate, $\lambda$ is the regularization parameter, $\boldsymbol{h}_{l}^{\mu}$ is the activation of layer $l$, and $\boldsymbol{\delta}_{l+1}^{\mu}$ is the error signal from the next layer.
By updating $P_{l}$ using the error signals, we ensure that the feedback weights of each layer are adapted to capture the most relevant directions in the error space, facilitating effective learning throughout the network. Notably, while we use the same learning rate $\eta$ and weight decay $\lambda$ for all components $\{W_{l}\}$, $\{Q_{l}\}$, and $\{P_{l}\}$, it can potentially differ.
In the interest of brevity, we omit simulations of deep linear networks, as the extension from the single-layer case is straightforward. Instead, we proceed directly to deep nonlinear networks, where the impact of constrained error feedback presents more complex and interesting dynamics.
4.   Nonlinear networks and complex architectures
Adapting our Restricted Adaptive Feedback (RAF) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable. The local update rules in eq. (11) remain the same; the primary difference lies in the introduction of nonlinear activation functions during the propagation of signals and errors. Specifically, the forward and backward passes are modified as follows:
$h_{l}=f(W_{l}h_{l-1}),\quad\delta_{l}=Q_{l}P_{l}\delta_{l+1}\odot f^{\prime}(h_{l}),$ (12)
where $f$ is the nonlinear activation function applied element-wise, and $f^{\prime}$ is its derivative.
To test whether our learning rule extends effectively from linear to nonlinear models, we trained deep networks on the CIFAR-10 dataset. We used a simple nonlinear model with four fully connected layers of 512 ReLU neurons each. While not state-of-the-art, this model provides a suitable testbed for evaluating our theory’s applicability to nonlinear architectures and complex data.
To isolate the impact of feedback dimensionality, we applied the RAF rule in eq. (11), constraining the rank of feedback matrices in one specific layer at a time while leaving the others unrestricted. We then measured the network’s test accuracy and compared it to backpropagation (BP) as a baseline. Figure 3a shows the accuracy as a function of the feedback rank $r_{l}$, with each curve representing a different constrained layer.
Consistent with our findings for the linear model, constraining the feedback rank to $r=d=10$, in any layer, match BP performance (Figure 3a), where $d$ is the number of classes. Interestingly, the shallower layers performed well even under tighter rank constraints ($r_{l}<d$), suggesting that the deeper layers compensate for the limited feedback in the earlier layers by effectively adjusting their weights.
This compensatory effect is possible because no information is lost during the feedforward pass, unlike in a bottlenecked network. To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of RAF-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b). The results confirm that RAF-trained networks leverage their width to maintain high performance despite feedback constraints.
Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using RAF can still match BP performance, demonstrating RAF’s robustness even with dimensional constraints across the entire network.
Task dimensionality determines the minimal rank
Our linear analysis suggests that the error signal dimensionality needed for effective learning is tied to the loss gradient dimensionality, which depends on the number of classes in the data. To test this, we trained networks on subsets of CIFAR-100 with 50, 75, and 100 classes, constraining the ranks of all feedback matrices to $r$ (Figure 3c).
The results show that network performance matches BP when the feedback rank equals the number of classes ($r=d$). It indicates that the minimal rank required for effective learning aligns with the task’s complexity, defined by the number of output classes.
Low-dimensional error signal from the top layer
Our theory shows how to propagate error signals from deeper layers to shallower ones using restricted adaptive feedback. However, error projections can also bypass intermediate layers entirely, leading to different variants of Direct Restricted Adaptive Feedback (dRAF), analogous to Direct Feedback Alignment [22]. For example, we can make direct connections from the output or the penultimate layer to earlier layers, training these connections using our algorithm (Figure 3d). As with RAF, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$. This model is particularly important because it is more flexible and has greater potential to explain learning in the brain, where error signals may arrive from different pathways.
Figure 3: Restricted Adaptive Feedback (RAF) efficiently trains nonlinear networks on CIFAR-10. (a) Constraining feedback in any layer to $r=d=10$ matches full BP performance. Different curves ($B_{1}$, $B_{2}$, and $B_{3}$) correspond to restricting feedback at different layers. (b) Reducing feedback dimensionality has minimal impact on performance, whereas reducing network width significantly degrades accuracy, indicating that high-dimensional representations are still utilized. (c) Minimal error dimensionality is sufficient for learning, as shown by subsampling classes from CIFAR-100. All feedback matrices are constrained to the same rank $r$. (d) Direct RAF: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance. (e) Performance of a 4-block VGG-like convolutional network trained with RAF, constraining the layers with 512 channels. Inset: Training the same network with all layers constrained to a fraction of their size; the smallest layer has 64 channels. (f) A transformer-based architecture trained to classify images, demonstrating that RAF generalizes to more advanced network architectures. See Supplemental Information for details on all implementations.
Convolutional neural networks
Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even on complex datasets, matching backpropagation (BP) performance. Here, we extend this investigation to convolutional architectures.
Training convolutional networks with Feedback Alignment (FA) is notoriously difficult [8, 23]. Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using Restricted Adaptive Feedback (RAF). We aim to determine whether convolutional networks can also benefit from low-dimensional error feedback.
We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see Appendix for details). Using RAF, we decoupled the error propagation from the feedforward pass in all layers. Initially, we constrained the feedback error only in the blocks containing 512 (Figure 3e). Consistent with our findings in fully connected networks, convolutional networks learn well with a feedback matrix with a rank similar to the number of classes, $d=10$.
To test whether convolutional networks can train when constraining all feedback paths, we further restricted each block to have feedback matrices with ranks equal to $1/2$, $1/4$, or $1/8$ of the block width (Figure 3e inset). Our results indicate that reducing the error dimensionality has a minimal impact on performance, except in the most extreme case. Specifically, constraining the feedback rank to $1/8$ of the block width resulted in a noticeable drop in performance. This finding aligns with our previous results, as the layer with 64 channels received feedback with a rank of $r=8$. Overall, our results demonstrate that convolutional networks can be efficiently trained using a minimal error signal.
Transformers
Transformers have become a dominant architecture in natural language processing and, more recently, have achieved state-of-the-art performance in image classification tasks [26, 27]. Unlike convolutional and standard feedforward architectures, transformers rely on a self-attention mechanism that computes interactions between tokens through the multiplication of key, query, and value matrices [28]. These operations introduce complex, nonlinear dependencies that make it unclear whether low-dimensional error feedback, as proposed in our framework, can effectively support learning in transformer models.
To investigate this question, we applied our approach to visual transformer architectures trained for image recognition on the CIFAR-10 dataset. Each weight matrix in the transformer architecture, including key, query, and value matrices, was trained using low-dimensional error feedback from the subsequent layer. Feedback matrices were trained using the RAF framework (see the appendix for details). The network’s performance as a function of the feedback rank, which was kept identical for all matrices, is shown in Figure 3f.
Our results indicate that, consistent with our findings in both nonlinear and convolutional networks, an error signal with dimensionality as low as the task dimensionality (i.e. the number of output classes) is sufficient to train transformer architectures effectively (Figure 3f). Despite the intricate nature of self-attention computations, training with low-dimensional feedback yielded performance comparable to conventional backpropagation on CIFAR-10 with rank $r=10$. Notably, training transformers with low-dimensional feedback using RAF not only matched but, in some cases, consistently outperformed backpropagation by a small margin. This improvement was observed even when hyperparameters were independently optimized for both BP and RAF. The improved generalization could be understood as a regularization effect achieved by constraining the feedback. However, since we observe it only in the transformer architecture, it is not a general property of a low-dimensional error signal, and we leave further investigation to a future study.
Figure 4: Receptive fields (RFs) are shaped by feedback dimensionality rather than feedforward constraints. (a) Model of the early ventral visual stream. Top: Retinal-to-cortical connectivity is modeled as a bottleneck in the network; adapted from [29]. Middle: A bottleneck is applied only to the feedback connections, restricting error dimensionality. Bottom: Using direct RAF (dRAF), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck. (b) Receptive fields (RFs) in a model trained with BP, corresponding to the setup in (a) top. Left: No bottleneck. Right: Bottleneck at the retinal output, which induces center-surround RFs. (c) Training the retinal output with RAF. Restricting the rank of the feedback matrix produces center-surround RFs. (d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dRAF. The presence of high-dimensional error signals results in oriented RFs.
5.   Error dimensionality shapes neural receptive fields
We have shown that neural networks can be efficiently trained using minimal error signals comparable to task dimensionality. Here, we investigate how error dimensionality affects neural representations, providing insights into receptive fields observed in the brain.
Lindsey et al. found that in convolutional models of the visual system, narrow feedforward bottlenecks between the retina and the brain led to center-surround receptive fields in the retinal layer, similar to mammals [29]. Wider bottlenecks resulted in orientation-selective receptive fields, as seen in salamanders. We hypothesize that these effects are due to constraints on the error signal reaching the retina, rather than the feedforward bottlenecks themselves.
To test this hypothesis, we trained a model similar to that of Lindsey et. al  [29] but with full-width layers throughout. Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with RAF (Figure 4c), thereby isolating the impact of error dimensionality on neural representations. We extracted the receptive fields of neurons in the retinal layer using visualization techniques [30] (Figure 4d).
Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer. To further validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dRAF to train the retinal layer without restricting the backward pathway. In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix).
This experiment demonstrates that error dimensionality influences neuronal tuning and neural representations. Specifically, lower-dimensional error signals promote higher symmetries in emergent receptive fields. Our findings underscore the importance of considering the dimensionality and pathways of error signals when studying neural computations in the brain.
6.   Discussion
Our work demonstrates that neural networks can be trained effectively using minimal error signals constrained to the task’s intrinsic dimensionality rather than the higher dimensionality of the network’s representations. By adopting a factorized version of Feedback Alignment with low-rank matrices and training both left and right spaces of the feedback matrix, we showed that deep networks—linear, nonlinear, and convolutional—can match the performance of full backpropagation even under stringent error-dimensionality constraints. This finding highlights that the essential information required for learning is tied to the complexity of the task, measured by the number of output classes, and does not scale with the number of parameters or size of the model. Additionally, we revealed that constraining error dimensionality influences neural representations, providing a potential explanation for biological phenomena, such as center-surround receptive fields in the retina.
This work aims to explore potential mechanisms for implementing gradient descent in the brain. Recognizing that high-dimensional feedback is not necessary for effective learning is a significant step toward developing more flexible and biologically realistic models of learning. This insight suggests that the brain might utilize low-dimensional error signals to drive learning processes, aligning with anatomical and physiological constraints.
Bridging biological constraints and gradient-based learning.
Our work aims to reconcile gradient-based learning with the anatomical constraints of cortical circuits. In the cortex, feedback connections are generally sparser, less structured, and more selective than their feedforward counterparts [31], suggesting that learning may rely on a restricted top-down error signal. Several alternative frameworks have been proposed to explain learning in hierarchical biological circuits. Predictive coding models suggest that the cortex continuously generates predictions about sensory input and computes discrepancies, or prediction errors, between expected and actual signals [32, 33] . Equilibrium propagation [34] offers another biologically inspired alternative, demonstrating how energy-based models can approximate backpropagation through local weight updates. While these approaches provide valuable insights into neural computation, they deviate from the conventional framework of gradient-based optimization. In contrast, our approach remains firmly within the gradient descent paradigm, allowing direct loss minimization. This key property enables its seamless application across diverse architectures, from fully connected nonlinear networks to convolutional models and transformer-based attention mechanisms.
Between reward-based learning and full gradient descent.
By showing that deep networks can be trained with low-dimensional error signals, our framework bridges the gap between gradient-based learning and reward-based learning. Studies have suggested that dopaminergic neurons in the ventral tegmental area (VTA) encode reward signals that capture multiple task attributes rather than a single scalar value [35, 36, 37]. Our findings provide a theoretical foundation for moving beyond scalar reward-based learning toward a framework where low-dimensional reward signals carry richer information about the underlying loss (or gain). This perspective positions our model as an intermediary between reinforcement learning’s scalar feedback and the full backpropagation of error gradients. More broadly, our results suggest that learning could emerge through the strategic compression of error signals into task-relevant representations, offering a biologically plausible alternative to high-dimensional gradient propagation.
Aligning to the error space.
On its surface, our novel learning rule is similar to previous Feedback Alignment (FA) schemes. However, it offers a conceptual novelty. In traditional FA, learning the feedback weights aims to align them with the feedforward weights [5, 10], mirroring full backpropagation. In contrast, by factorizing the feedback matrix $B=QP$, we also align the row space of the feedback matrix with the source of the error, thereby improving the quality of the error signal itself.
Computational complexity.
Despite the additional requirement of training the feedback matrices, the computational cost of using low-rank feedback does not increase significantly. The updates for the low-rank matrices are linear in the number of neurons or filters per layer, while the forward pass and the primary weight updates remain quadratic [38]. This efficiency makes our approach practical for large-scale networks and suggests that similar mechanisms could be feasible in biological neural networks.
Gradient descent in high dimensions.
Our findings invite a rethinking of gradient descent dynamics in large, overparameterized networks. Typically, the weight dynamics during training are high-dimensional [39]. However, when the error signal is low-dimensional, the weight updates in each layer are confined to a much lower-dimensional subspace. This constraint could have implications for understanding the generalization capabilities of neural networks, as it suggests that effective learning does not require exploring the full parameter space. Indeed, previous studies found that the Hessian of the loss function shows low-rank structure [40]. Exploring this connection could open new avenues for understanding the dynamics of gradient descent in high-dimensional loss spaces [41].
Exploring the potential of low-dimensional feedback in more complex architectures.
This approach not only improves learning efficiency but may also act as a form of regularization, as can be seen from the improved performance of low-rank feedback matrices on visual transformers—the possibility of using restricted error signals as a form of regularization warrants further investigation.
In summary, our findings not only challenge the conventional reliance on high-dimensional error propagation but also pave the way for the development of biologically inspired, computationally efficient learning algorithms that hold promise for advancing both artificial intelligence and our understanding of brain function." This would leave the reader with a clear impression of both the scientific and translational impact of your work
Acknowledgments
This work was supported by the Gatsby Charitable Foundation and the Azrieli Foundation.
Appendix
Appendix A Full linear theory
in this section, we provide the analysis of the linear networks for both Feedback Alignment and RAF
Detailed analysis of Feedback Alignment learning dynamics
From eq. (4) in the main text, the weight updates for a single sample $\mu$ are given by:
$\Delta W_{1}^{\mu}$ $=\eta B\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right)\boldsymbol{x}^{\mu\top},$ (A1)
$\Delta W_{2}^{\mu}$ $=\eta\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right)\boldsymbol{x}^{\mu\top}W_{1}^{\top},$
where:
  • $\eta$ is the learning rate,
  • $\boldsymbol{x}^{\mu}\in\mathbb{R}^{n}$ is the input vector for sample $\mu$,
  • $\boldsymbol{y}^{\mu}\in\mathbb{R}^{m}$ is the corresponding target output,
  • $W_{1}\in\mathbb{R}^{k\times n}$ and $W_{2}\in\mathbb{R}^{m\times k}$ are the weight matrices,
  • $B\in\mathbb{R}^{k\times m}$ is a predefined matrix (e.g., a feedback or scaling matrix).
We introduce the empirical covariance matrices:
$\Sigma_{io}$ $=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}\boldsymbol{x}^{\mu\top},$ (A2)
$\Sigma_{oo}$ $=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}\boldsymbol{y}^{\mu\top},$ (A3)
$\Sigma_{ii}$ $=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{x}^{\mu}\boldsymbol{x}^{\mu\top}=I,$ (A4)
where $\Sigma_{ii}=I$ assumes that the input vectors are whitened (i.e., have unit covariance). Summing over all $p$ training examples, we obtain the average weight updates:
$\Delta W_{1}$ $=\frac{\eta}{p}\sum_{\mu=1}^{p}\Delta W_{1}^{\mu}$ (A5)
$=\frac{\eta}{p}\sum_{\mu=1}^{p}B\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right)\boldsymbol{x}^{\mu\top}.$
Using these definitions, the update for $W_{1}$ simplifies to:
$\Delta W_{1}$ $=\eta B\left(\Sigma_{io}-W_{2}W_{1}\Sigma_{ii}\right)$ (A6)
$=\eta B\left(\Sigma_{io}-W_{2}W_{1}\right).$ (A7)
Under the limit as $\eta\rightarrow 0$ with $\eta=\frac{dt}{\tau}$ (where $\tau$ is a time constant), we transition from discrete updates to continuous-time dynamics:
$\tau\frac{dW_{1}}{dt}=B\left(\Sigma_{io}-W_{2}W_{1}\right),$ (A8)
which matches the weight dynamics presented in eq. (5) of the main text.
Similarly, the update for $W_{2}$ becomes:
$\Delta W_{2}$ $=\frac{\eta}{p}\sum_{\mu=1}^{p}\Delta W_{2}^{\mu}$ (A9)
$=\frac{\eta}{p}\sum_{\mu=1}^{p}\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right)\boldsymbol{x}^{\mu\top}W_{1}^{\top}$ (A10)
$=\eta\left(\Sigma_{io}-W_{2}W_{1}\right)W_{1}^{\top},$ (A11)
which simplifies under the same limit to:
$\tau\frac{dW_{2}}{dt}=\left(\Sigma_{io}-W_{2}W_{1}\right)W_{1}^{\top}.$ (A12)
Eq. (A8) and eq. (A12) describe the continuous-time dynamics of the weights $W_{1}$ and $W_{2}$ under the given learning rule.
We consider the singular value decomposition (SVD) of the covariance matrix $\Sigma_{io}$:
$\Sigma_{io}=USV^{\top},$ (A13)
where:
  • $U\in\mathbb{R}^{m\times d}$ and $V\in\mathbb{R}^{n\times d}$ are matrices with orthonormal columns,
  • $S\in\mathbb{R}^{d\times d}$ is a diagonal matrix containing the singular values,
  • $d$ is the rank of $\Sigma_{io}$.
We perform a rotation of the weight matrices and $B$ as follows:
$W_{1}$ $=\bar{W}_{1}V^{\top},$ (A14)
$W_{2}$ $=U\bar{W}_{2},$
$B$ $=\bar{B}U^{\top}.$
Substituting these into the previous weight dynamics, we have for $W_{1}$:
$\tau\frac{dW_{1}}{dt}$ $=B\left(\Sigma_{io}-W_{2}W_{1}\right)$ (A15)
$=\bar{B}U^{\top}\left(USV^{\top}-U\bar{W}_{2}\bar{W}_{1}V^{\top}\right).$ (A16)
Since $U^{\top}U=I$ (due to orthonormal columns of $U$), we can simplify:
$\tau\frac{dW_{1}}{dt}$ $=\bar{B}\left(U^{\top}U\right)\left(S-\bar{W}_{2}\bar{W}_{1}\right)V^{\top}$ (A17)
$=\bar{B}\left(S-\bar{W}_{2}\bar{W}_{1}\right)V^{\top}.$ (A18)
Recognizing that $W_{1}=\bar{W}_{1}V^{\top}$, we can write $\frac{dW_{1}}{dt}=\frac{d\bar{W}_{1}}{dt}V^{\top}$. Thus, multiplying both sides on the right by $V$ (since $V^{\top}V=I$):
$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}\left(S-\bar{W}_{2}\bar{W}_{1}\right).$ (A19)
Similarly, for $W_{2}$:
$\tau\frac{dW_{2}}{dt}$ $=\left(\Sigma_{io}-W_{2}W_{1}\right)W_{1}^{\top}$ (A20)
$=\left(USV\top-U\bar{W}_{2}\bar{W}_{1}V^{\top}\right)\left(\bar{W}_{1}V^{\top}\right)^{\top}$ (A21)
Using the fact that $V^{\top}V=I$ we simplify:
$\tau\frac{dW_{2}}{dt}$ $=\left(US-U\bar{W}_{2}\bar{W}_{1}\right)\bar{W}_{1}^{\top}$ (A23)
Since $W_{2}=U\bar{W}_{2}$, we have $\frac{dW_{2}}{dt}=U\frac{d\bar{W}_{2}}{dt}$. Multiplying both sides on the left by $U^{\top}$:
$\tau\frac{d\bar{W}_{2}}{dt}=\left(S-\bar{W}_{2}\bar{W}_{1}\right)\bar{W}_{1}^{\top}.$ (A25)
Eqs. (A19) and (A25) describe the dynamics of the rotated weights $\bar{W}_{1}$ and $\bar{W}_{2}$:
$\tau\frac{d\bar{W}_{1}}{dt}$ $=\bar{B}\left(S-\bar{W}_{2}\bar{W}_{1}\right),$ (A26)
$\tau\frac{d\bar{W}_{2}}{dt}$ $=\left(S-\bar{W}_{2}\bar{W}_{1}\right)\bar{W}_{1}^{\top}.$
Full derivation of Restricted Adaptive Feedback (RAF) learning dynamics
We consider an alternative algorithm where the matrix $B$ is replaced by $B=QP$, with $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$.
The updates for $Q$ and $P$ for a single sample $\mu$ are given by:
$\Delta Q^{\mu}$ $=\eta W_{1}\boldsymbol{x}^{\mu}P\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right),$ (A27)
$\Delta P^{\mu}$ $=\eta P\boldsymbol{y}^{\mu}\boldsymbol{y}^{\mu\top}\left(I-P^{\top}P\right).$
Summing over all $p$ training examples, we obtain the average updates:
$\Delta Q$ $=\frac{\eta}{p}\sum_{\mu=1}^{p}\Delta Q^{\mu}=\eta W_{1}\left(\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{x}^{\mu}\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right)^{\top}P^{\top}\right),$ (A28)
$\Delta P$ $=\frac{\eta}{p}\sum_{\mu=1}^{p}\Delta P^{\mu}=\eta P\left(\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}\boldsymbol{y}^{\mu\top}\right)\left(I-P^{\top}P\right).$
We simplify the update for $Q$:
$\Delta Q$ $=\eta W_{1}\left(\left(\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{x}^{\mu}\left(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}\right)^{\top}\right)P^{\top}\right)$ (A29)
$=\eta W_{1}\left(\left(\Sigma_{io}^{\top}-W_{1}^{\top}W_{2}^{\top}\Sigma_{ii}\right)P^{\top}\right)$ (A30)
$=\eta W_{1}\left(\left(\Sigma_{io}-W_{2}W_{1}\right)^{\top}P^{\top}\right)$ (A31)
Similarly, the update for $P$ simplifies to:
$\Delta P$ $=\eta P\Sigma_{oo}\left(I-P^{\top}P\right).$ (A32)
Thus, the updates for $Q$ and $P$ become:
$\Delta Q$ $=\eta W_{1}\left(\Sigma_{io}-W_{2}W_{1}\right)^{\top}P^{\top},$ (A33)
$\Delta P$ $=\eta P\Sigma_{oo}\left(I-P^{\top}P\right).$
Under the continuous-time assumption, where $\eta\rightarrow 0$ with $\eta=\frac{dt}{\tau}$, the updates for $Q$ and $P$ become differential equations:
$\tau\frac{dQ}{dt}=W_{1}\left(\Sigma_{io}-W_{2}W_{1}\right)^{\top}P^{\top},$ (A34)
and
$\tau\frac{dP}{dt}=P\Sigma_{oo}\left(I-P^{\top}P\right).$ (A35)
Finally, substituting $B=QP$ into the update for $W_{1}$ from eq. A8, we find that the dynamics for $W_{1}$ become:
$\tau\frac{dW_{1}}{dt}=QP\left(\Sigma_{io}-W_{2}W_{1}\right)$ (A36)
We perform rotations similar to before:
$W_{1}$ $=\bar{W}_{1}V^{\top},$ (A37)
$W_{2}$ $=U\bar{W}_{2},$
$P$ $=\bar{P}U^{\top}.$
Substituting for $Q$ we get:
$\tau\frac{dQ}{dt}=\bar{W}_{1}V_{1}^{\top}\left(USV^{\top}-U\bar{W}_{2}\bar{W}_{1}V^{\top}\right)^{\top}(PU^{\top})^{\top}$ (A38)
$=\bar{W}_{1}V_{1}^{\top}V\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{T}U^{\top}U\bar{P}^{\top}$ (A39)
$=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$ (A40)
Substituting these rotations into the previous derivations, we update the dynamics.
For $W_{1}$, the update equation is:
$\tau\frac{dW_{1}}{dt}=B\left(\Sigma_{io}-W_{2}W_{1}\right).$ (A41)
Since $B=QP=Q\bar{P}U^{\top}$, $W_{1}=\bar{W}_{1}V^{\top}$, $W_{2}=U\bar{W}_{2}$, and $\Sigma_{io}=USV^{\top}$, we have:
$\tau\frac{d(\bar{W}_{1}V^{\top})}{dt}$ $=Q\bar{P}U^{\top}\left(USV^{\top}-U\bar{W}_{2}\bar{W}_{1}V^{\top}\right)$ (A42)
$=Q\bar{P}\left(SV^{\top}-\bar{W}_{2}\bar{W}_{1}V^{\top}\right).$ (A43)
Since $V^{\top}$ is constant, we can write:
$\tau\frac{d\bar{W}_{1}}{dt}V^{\top}=Q\bar{P}\left(SV^{\top}-\bar{W}_{2}\bar{W}_{1}V^{\top}\right).$ (A44)
Multiplying both sides on the right by $V$ (using $V^{\top}V=I$):
$\tau\frac{d\bar{W}_{1}}{dt}=Q\bar{P}\left(S-\bar{W}_{2}\bar{W}_{1}\right).$ (A45)
Substituting $P=\bar{P}U^{\top}$ and $\Sigma_{oo}=US^{2}U^{\top}$ in (A35), we get:
$\tau\frac{d(\bar{P}U^{\top})}{dt}$ $=\bar{P}U^{\top}\left(US^{2}U^{\top}\right)\left(I-U\bar{P}^{\top}\bar{P}U^{\top}\right)$ (A46)
$=\bar{P}S^{2}\left(I-\bar{P}^{\top}\bar{P}\right)U^{\top}.$ (A47)
Multiplying both sides on the right by $U$ (since $U^{\top}U=I$):
$\tau\frac{d\bar{P}}{dt}=\bar{P}S^{2}\left(I-\bar{P}^{\top}\bar{P}\right).$ (A48)
In summary, under the rotations, the updated dynamics are:
$\tau\frac{d\bar{W}_{1}}{dt}$ $=Q\bar{P}\left(S-\bar{W}_{2}\bar{W}_{1}\right),$ (A49)
$\tau\frac{d\bar{W}_{2}}{dt}$ $=\left(S-\bar{W}_{2}\bar{W}_{1}\right)\bar{W}_{1}^{\top},$
$\tau\frac{d\bar{P}}{dt}$ $=\bar{P}S^{2}\left(I-\bar{P}^{\top}\bar{P}\right).$
$\tau\frac{d\bar{Q}}{dt}$ $=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$
Appendix B Learning rule implementation
Our implementation and optimization were conducted entirely using PyTorch. For the RAF layers, we initialized all weights using Kaiming uniform initialization. We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule. In the output layer, we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$. In the hidden layers, $\boldsymbol{P}$ is learned to project onto the principal directions of the error signal from the subsequent layer, represented as $\boldsymbol{\delta}_{l+1}$. Specifically, for a layer $l$ with input dimension $n$, output dimension $m$, and rank constraint $r$, we proceed as follows:
Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$, $P_{l}\in\mathbb{R}^{r\times m}$, weight matrix $W_{l}\in\mathbb{R}^{m\times n}$
Output: Gradient w.r.t. input grad_input, updates $\Delta Q_{l}$, $\Delta P_{l}$, $\Delta W_{l}$
Compute covariance matrix:
if use_targets then
        $C\leftarrow(\boldsymbol{y}-\boldsymbol{\bar{y}})^{\top}(\boldsymbol{y}-\boldsymbol{\bar{y}})$;
else
        $C\leftarrow(\boldsymbol{\delta_{l+1}}-\boldsymbol{\bar{\delta_{l+1}}})^{\top}(\boldsymbol{\delta_{l+1}}-\boldsymbol{\bar{\delta_{l+1}}})$;
$\gamma\leftarrow\displaystyle\max\left(\operatorname{diag}\left(C\right)\right)$;
Compute ’gradient’ w.r.t. input:
[…]
Compute updates for $Q_{l}$ and $P_{l}$:
$\Delta Q_{l}\leftarrow\boldsymbol{h}_{l}^{\top}\left(P_{l}\boldsymbol{\delta}_{l+1}^{\top}\right)^{\top}$;
$\Delta P_{l}\leftarrow P_{l}\left(\dfrac{C}{\gamma}\right)\left(I-P_{l}^{\top}P_{l}\right)$;
Compute update for $W_{l}$:
$\Delta W_{l}\leftarrow\boldsymbol{\delta}_{l+1}^{\top}\boldsymbol{h}_{l}$;
Algorithm 1 Modified Backward Pass for RAF Layer
The derivation of the dRAF (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
Appendix C Numerical experiments
Layer-Wise Constraints
In Figure [3.a], we present the results of training a network with four hidden layers, each containing 512 neurons, and an output layer with 10 neurons on the CIFAR-10 dataset. The network was trained using RAF, as described earlier, without rank constraints, except for one layer at a time. For comparison, we also trained the network using standard backpropagation as a baseline. All networks were trained with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975. Each experiment was repeated 10 times.
Constraining All Layers
For the results shown in Figure [3.b], we trained the same network with four hidden layers, each containing 512 neurons. This time, we applied rank constraints to all layers simultaneously, with rank values $r=64,32,16,10$.
To further demonstrate that the network still utilizes high-rank representations, we also trained a variant of the network with 64 neurons in each hidden layer, without applying any rank constraints. All training was conducted with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training carried out for 160 epochs and an exponential learning rate decay factor of 0.975. Each experiment was repeated 10 times.
CIFAR-100 Sub-sampling
For the results shown in Figure [3.c], we trained the same model on the CIFAR-100 dataset, sampling different numbers of classes $d$, with $d=50,75,100$. For each sub-sample, we trained the model while applying rank constraints to all layers, using various rank values. This was done to demonstrate that the dimensionality of the error signal depends on the task dimensionality $d$. All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975. Each training run was repeated 5 times.
dRAF
For the results shown in Figure [3.d], we trained the same model using dRAF. In one experiment, we propagated the error signal from the last layer to all preceding layers. In a separate experiment, we propagated the error signal directly from the penultimate layer to all earlier layers, applying rank constraints to these layers while keeping the last layer at full rank. All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, training for 160 epochs with an exponential learning rate decay factor of 0.975. Each training run was repeated 5 times
Convolutional Neural Networks
To extend our RAF algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal. This effectively constrains the dimensionality of the error signal across the spatial dimensions. Similar to the implementation for fully connected layers in Algorithm 1, we modify the backward pass in the same way for convolutional layers.
The key difference in the convolutional context is how the projection matrix $\boldsymbol{P}$ operates on the error signals. In convolutional layers, $\boldsymbol{P}$ functions as a $1\times 1$ convolutional filter, projecting the error signal at each spatial location from $m$ channels down to $r$ channels. This reduces the error signal’s dimensionality to $r$ per pixel, adhering to the rank constraint.
As in the fully connected case, the matrix $\boldsymbol{P}$ is learned using Oja’s rule. However, the covariance matrix $\boldsymbol{C}$ is computed over both the batch and spatial dimensions—that is, across all pixels in all images within the batch. This approach captures the covariance structure of pixel representations more effectively, enabling $\boldsymbol{P}$ to project the error signals appropriately in the convolutional setting.
VGG-like Architecture
We trained a VGG-like convolutional neural network comprising four convolutional blocks, as detailed in Table C1. For the experiments presented in Figure 4(a), a rank constraint was applied exclusively to the final convolutional block, which consists of 512 channels. In contrast, for the results shown in Figure 4(b), rank constraints were imposed on all layers, reducing the rank to 1/2, 1/4, and 1/8 of the original channel dimensions.
Each network was trained using the Adam optimizer with a learning rate of $5\times 10^{-4}$, a weight decay of $5\times 10^{-5}$, and an exponential learning rate decay factor of 0.98. Training was conducted for 250 epochs, with each experiment repeated five times to ensure robustness.
Vision Transformer (ViT)
We implemented a Vision Transformer (ViT) model that consists of seven self-attention blocks, each with an embedding dimension of 384 and four attention heads. The input images were partitioned into nonoverlapping patches, which were linearly projected into a 384-dimensional feature space. These patch embeddings were augmented with learned positional encodings and processed through a stack of seven transformer encoder layers. Each encoder block employed multi-head self-attention with four heads, followed by pre-norm layer normalization and dropout for regularization. The final representation was obtained from a learned class token, which was subsequently processed by a fully connected layer for classification.
Training was carried out using a batch size of 32, a learning rate of $3\times 10^{-4}$, and a weight loss of $1\times 10^{-4}$. We used the Adam optimizer with AMSGrad, training for 150 epochs with an exponential learning rate decay factor of 0.98. Each experiment was repeated five times to ensure robustness.
Neural Receptive Fields
To analyze neural receptive fields, we trained the same convolutional network architecture as used in Lindsey et. al. [29]. The model consists of two convolutional layers with ReLU activations to simulate retinal processing, followed by three additional convolutional layers with ReLU activations to model the ventral visual stream (VVS). Fully connected layers were used for classification, with the complete architecture detailed in Table C2. A bottleneck was introduced between the retinal and VVS layers to constrain the information flow.
Training was performed using the Restricted Adaptive Feedback (RAF) method, where rank constraints were applied to feedback sent to the retina with ranks $r=2,4,32$. The network was trained with hyperparameters similar to those in Linsdey et. al. [29], using the RMSProp optimizer with a learning rate of $1\times 10^{-4}$, a weight decay of $1\times 10^{-5}$, and an exponential learning rate decay factor of 0.985. Each model was trained for 120 epochs, and all experiments were repeated five times to ensure robustness.
Additionally, we conducted experiments with dynamic Restricted Adaptive Feedback (dRAF), where feedback to the retina originated from higher visual layers while feedforward connections within the retina were constrained to four channels.
Table C1: CNN Architecture
Layer(s) Output Size Details
Input $32\times 32\times 3$
Convolutional Block 1
Conv2D + ReLU $32\times 32\times 64$ $3\times 3$ conv, 64 filters, padding=1
BatchNorm2D $32\times 32\times 64$
Conv2D + ReLU $32\times 32\times 64$ $3\times 3$ conv, 64 filters, padding=1
BatchNorm2D $32\times 32\times 64$
MaxPool2D $16\times 16\times 64$ $2\times 2$ max pool, stride=2
Convolutional Block 2
Conv2D + ReLU $16\times 16\times 128$ $3\times 3$ conv, 128 filters, padding=1
BatchNorm2D $16\times 16\times 128$
Conv2D + ReLU $16\times 16\times 128$ $3\times 3$ conv, 128 filters, padding=1
BatchNorm2D $16\times 16\times 128$
MaxPool2D $8\times 8\times 128$ $2\times 2$ max pool, stride=2
Convolutional Block 3
Conv2D + ReLU $8\times 8\times 256$ $3\times 3$ conv, 256 filters, padding=1
BatchNorm2D $8\times 8\times 256$
Conv2D + ReLU $8\times 8\times 256$ $3\times 3$ conv, 256 filters, padding=1
BatchNorm2D $8\times 8\times 256$
MaxPool2D $4\times 4\times 256$ $2\times 2$ max pool, stride=2
Convolutional Block 4
Conv2D + ReLU $4\times 4\times 512$ $3\times 3$ conv, 512 filters, padding=1
BatchNorm2D $4\times 4\times 512$
Conv2D + ReLU $4\times 4\times 512$ $3\times 3$ conv, 512 filters, padding=1
BatchNorm2D $4\times 4\times 512$
AdaptiveAvgPool2D $1\times 1\times 512$ Output size $(1,1)$
Flatten $512$ Flatten to vector
Classifier
Fully Connected + ReLU $256$ Linear layer, $512\rightarrow 256$
Dropout $256$ Dropout probability $p=0.4$
Fully Connected $C$ Linear layer, $256\rightarrow C$
Table C2: Retina model Architecture
Layer(s) Output Size Details
Input $32\times 32\times 1$ Grayscale input
Retina
Conv2D + ReLU $32\times 32\times 32$ $9\times 9$ conv, 32 filters, padding=4
Conv2D + ReLU $32\times 32\times 32$ $9\times 9$ conv, 32 filters, padding=4
VVS
Conv2D + ReLU $32\times 32\times 32$ $9\times 9$ conv, 32 filters, padding=4
Conv2D + ReLU $32\times 32\times 32$ $9\times 9$ conv, 32 filters, padding=4
Conv2D + ReLU $32\times 32\times 32$ $9\times 9$ conv, 32 filters, padding=4
Flatten $32,768$ Flatten to vector
Classifier
Fully Connected + ReLU $1,024$ Linear layer, $32,768\rightarrow 1,024$
Dropout $1,024$ Dropout probability $p=0.5$
Fully Connected $C$ Linear layer, $1,024\rightarrow C$

### v2

@@
 Ultimately, our work offers a fresh perspective on neural network optimization and contributes to understanding learning mechanisms in both artificial and biological systems.
+Significance statement
+Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron, a process with limited biological plausibility.
+Our work demonstrates that neural networks can be trained effectively using low-dimensional error signals, comparable to the inherent complexity of the task rather than the size of the network.
+By developing a novel, biologically plausible learning paradigm based on adaptive low-dimensional feedback, we show theoretically and empirically that minimal error dimensionality can achieve competitive performance even in complex neural architectures.
+These findings challenge existing assumptions about neural optimization and suggest fundamental constraints that may shape learning in both artificial and biological systems.
 1.   Introduction
@@
 Our method exploits the low-dimensional nature of real-world tasks to train large networks efficiently without compromising their representational capacity, achieving performance on par with backpropagation.
-Our theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance.
-Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is both feasible and universal.
+A theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance.
+Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is feasible and universal.
 Finally, we explore how error dimensionality influences neural representations by examining its effects on receptive fields in an early visual system model, underscoring the broader implications of our findings.
@@
 However, no previous work has considered training with a constrained error pathway, and the effects of error dimensionality and training have not been systematically studied.
-In this work, our aim is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of neural networks.
-To this end, we introduce a novel learning scheme, Restricted Adaptive Feedback (RAF), that allows flexible control over the dimensionality of the errors (Figure 1).
+In this work, our objective is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of large neural networks.
+To this end, we introduce a novel learning scheme that enables deep layers to be learned using low-rank feedback matrices, which allow flexible control over the dimensionality of the loss gradients in each layer (Figure 1).
 Our main contributions are:
   1. 1.
-    We present a novel learning rule, Restricted Adaptive Feedback (RAF), which matches BP performance while requiring minimal error signals.
-We provide a detailed derivation of the learning dynamics in a simple linear case, establishing a foundational understanding of how RAF operates.
+    We present a novel local learning rule for low-rank feedback weights, which matches BP performance while requiring minimal error signals.
+We provide a detailed derivation of the learning dynamics in a simple linear case, establishing the foundation for why and how the feedback weights should be learned.
   2. 2.
-    We demonstrate that nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals, highlighting the versatility of RAF in practical scenarios.
+    We demonstrate that large and deep nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals and highlight that the feedback dimensionality should scale with the task rather than the size of the network.
   3. 3.
-    We show that more complex yet highly useful architectures, such as convolutional networks and transformers, can also be effectively trained with low-dimensional feedback.
+    We successfully train complex yet highly useful architectures, such as convolutional networks and visual transformers, using low-dimensional feedback.
   4. 4.
-    We reveal that error dimensionality shapes the receptive fields in a model of the ventral visual system, offering new insights into the relationship between learning mechanisms and biological neural representations.
+    We study a simplified model of the ventral visual system and show that the emerging receptive fields depend on the error-feedback dimensionality.
+These findings offer new insights into the relationship between learning mechanisms and biological neural representations.
 In the final section, we discuss the broader implications of our results for both neuroscience and machine learning.
@@
 Feedback Alignment (FA) replaces the transposed weights with fixed random feedback matrices, which gradually align with the forward weights during training.
-Restricted Adaptive Feedback (RAF) constrains error propagation through low-rank feedback matrices, limiting error dimensionality and preventing an exact mirroring of BP.
-Direct Restricted Adaptive Feedback (dRAF) extends RAF by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
+Our work framework provides low-dimensional feedback (LDF) using low-rank matrices, allowing full control of the dimnsionality of the gradients in each layer.
+Finally, direct low-dimensional feedback (dLDF) extends our LDF architecture by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
 3.   Low-dimensional error feedback in linear networks
@@
 The goal is to learn the low-dimensional structure of $A$ from the $p$ samples using a linear neural network.
-For simplicity, we assume that $p$ is sufficiently large to allow the network to fully recover the structure of $A$.
+For simplicity, we assume that $p$ is large enough to allow the network to fully recover the structure of $A$.
 A linear network model.
@@
 For a given data point $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}$, the weight updates, derived from the FA framework, are given by:
-$\Delta W_{1}^{\mu}$ $=\eta B^{T}(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
+$\Delta W_{1}^{\mu}$ $=\eta B(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
 $\Delta W_{2}^{\mu}$ $=\eta(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T}W_{1}^{T}.$
@@
 Assuming a small learning rate $\eta\ll 1$ with full-batch updates, we express the weight updates in continuous time:
-$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}^{T}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
+$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
 where $\eta=dt/\tau$.
@@
 The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$.
-(b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$, with $r=64$ and $r=8$, respectively.
+(b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=64$ and $r=8$, respectively.
 The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$.
-(d, e) Same as (b, c) but for Restricted Adaptive Feedback (RAF), where both $Q$ and $P$ are trained.
+(d, e) Same as (b, c) but training both and $Q$ and $P$ using Eqs (7) and (8) .
 (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully.
-(g) Training $P$ to align with the error, as in RAF, ensures full recovery of the top $r$ singular components, improving convergence and alignment.
+(g) Training $P$ to align with the error ensures full recovery of the top $r$ singular components, improving convergence and alignment.
 Stationary solutions for the training.
@@
 Taking the continuous-time limit (with $\eta\to 0$ and $\eta p=\tau$), we obtain a set of differential equations that describe the learning dynamics of the forward and backward weights:
-$\tau\frac{d\bar{W}_{1}}{dt}$ $=\bar{B}^{T}(S-\bar{W}_{2}\bar{W}_{1}),$ (9)
+$\tau\frac{d\bar{W}_{1}}{dt}$ $=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),$ (9)
 $\tau\frac{d\bar{W}_{2}}{dt}$ $=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T}-\lambda\bar{W}_{2},$
@@
 Notably, the updates to the feedback weights are local and follow learning rules that were well-studied in theoretical neuroscience [18, 19, 20, 21].
-Local plasticity makes our framework an attractive alternative for backpropagation in models of brain circuits.
-We refer to this learning framework as Restricted Adaptive Feedback (RAF).
-Figure 2 compares the learning dynamics of RAF with those of BP and FA where only the $Q$ matrix is learned.
-The results demonstrate that RAF effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality.
-Furthermore, RAF will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
+Local plasticity makes our framework an attractive alternative for backpropagation in brain circuit models.
+We refer to this learning framework as low-dimensional feedback alignment (LDFA).
+Figure 2 compares the learning dynamics of LDFA with those of BP and FA where only the $Q$ matrix is learned.
+The results demonstrate that LDFA effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality.
+Furthermore, LDFA will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
 3.3 Restricted adaptive feedback in deep architectures
@@
 4.   Nonlinear networks and complex architectures
-Adapting our Restricted Adaptive Feedback (RAF) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable.
+Adapting our low-dimensional feedback alignment (LDFA) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable.
 The local update rules in eq.
@@
 While not state-of-the-art, this model provides a suitable testbed for evaluating our theory’s applicability to nonlinear architectures and complex data.
-To isolate the impact of feedback dimensionality, we applied the RAF rule in eq.
+To isolate the impact of feedback dimensionality, we applied the LDFA rule in eq.
 (11), constraining the rank of feedback matrices in one specific layer at a time while leaving the others unrestricted.
@@
 This compensatory effect is possible because no information is lost during the feedforward pass, unlike in a bottlenecked network.
-To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of RAF-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b).
-The results confirm that RAF-trained networks leverage their width to maintain high performance despite feedback constraints.
-Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using RAF can still match BP performance, demonstrating RAF’s robustness even with dimensional constraints across the entire network.
+To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of LDFA-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b).
+The results confirm that LDFA-trained networks leverage their width to maintain high performance despite feedback constraints.
+Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using LDFA can still match BP performance, demonstrating LDFA’s robustness even with dimensional constraints across the entire network.
 Task dimensionality determines the minimal rank
@@
 Our theory shows how to propagate error signals from deeper layers to shallower ones using restricted adaptive feedback.
-However, error projections can also bypass intermediate layers entirely, leading to different variants of Direct Restricted Adaptive Feedback (dRAF), analogous to Direct Feedback Alignment [22].
+However, error projections can also bypass intermediate layers entirely, leading to different variants of direct low-dimensional feedbackalignment (dLDFA), analogous to Direct Feedback Alignment [22].
 For example, we can make direct connections from the output or the penultimate layer to earlier layers, training these connections using our algorithm (Figure 3d).
-As with RAF, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$.
+As with LDFA, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$.
 This model is particularly important because it is more flexible and has greater potential to explain learning in the brain, where error signals may arrive from different pathways.
-Figure 3: Restricted Adaptive Feedback (RAF) efficiently trains nonlinear networks on CIFAR-10.
+Figure 3: Low-dimensional feedback alignment (LDFA) efficiently trains nonlinear networks on CIFAR-10.
 (a) Constraining feedback in any layer to $r=d=10$ matches full BP performance.
@@
 All feedback matrices are constrained to the same rank $r$.
-(d) Direct RAF: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance.
-(e) Performance of a 4-block VGG-like convolutional network trained with RAF, constraining the layers with 512 channels.
+(d) Direct LDFA: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance.
+(e) Performance of a 4-block VGG-like convolutional network trained with LDFA, constraining the layers with 512 channels.
 Inset: Training the same network with all layers constrained to a fraction of their size; the smallest layer has 64 channels.
-(f) A transformer-based architecture trained to classify images, demonstrating that RAF generalizes to more advanced network architectures.
+(f) A transformer-based architecture trained to classify images, demonstrating that LDFA generalizes to more advanced network architectures.
 See Supplemental Information for details on all implementations.
 Convolutional neural networks
-Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even on complex datasets, matching backpropagation (BP) performance.
+Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even with complex datasets, matching backpropagation (BP) performance.
 Here, we extend this investigation to convolutional architectures.
 Training convolutional networks with Feedback Alignment (FA) is notoriously difficult [8, 23].
-Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using Restricted Adaptive Feedback (RAF).
+Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using LDFA.
 We aim to determine whether convolutional networks can also benefit from low-dimensional error feedback.
-We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see Appendix for details).
-Using RAF, we decoupled the error propagation from the feedforward pass in all layers.
+We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see the appendix for details).
+Using LDFA, we decoupled error propagation from the feedforward pass in all layers.
 Initially, we constrained the feedback error only in the blocks containing 512 (Figure 3e).
@@
 Each weight matrix in the transformer architecture, including key, query, and value matrices, was trained using low-dimensional error feedback from the subsequent layer.
-Feedback matrices were trained using the RAF framework (see the appendix for details).
+Feedback matrices were trained using the LDFA framework (see the appendix for details).
 The network’s performance as a function of the feedback rank, which was kept identical for all matrices, is shown in Figure 3f.
@@
 Despite the intricate nature of self-attention computations, training with low-dimensional feedback yielded performance comparable to conventional backpropagation on CIFAR-10 with rank $r=10$.
-Notably, training transformers with low-dimensional feedback using RAF not only matched but, in some cases, consistently outperformed backpropagation by a small margin.
-This improvement was observed even when hyperparameters were independently optimized for both BP and RAF.
+Notably, training transformers with low-dimensional feedback using LDFA not only matched but, in some cases, consistently outperformed backpropagation by a small margin.
+This improvement was observed even when hyperparameters were independently optimized for both BP and LDFA.
 The improved generalization could be understood as a regularization effect achieved by constraining the feedback.
@@
 Middle: A bottleneck is applied only to the feedback connections, restricting error dimensionality.
-Bottom: Using direct RAF (dRAF), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck.
+Bottom: Using direct LDFA (dLDFA), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck.
 (b) Receptive fields (RFs) in a model trained with BP, corresponding to the setup in (a) top.
@@
 Right: Bottleneck at the retinal output, which induces center-surround RFs.
-(c) Training the retinal output with RAF.
+(c) Training the retinal output with LDFA.
 Restricting the rank of the feedback matrix produces center-surround RFs.
-(d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dRAF.
+(d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dLDFA.
 The presence of high-dimensional error signals results in oriented RFs.
@@
 Wider bottlenecks resulted in orientation-selective receptive fields, as seen in salamanders.
-We hypothesize that these effects are due to constraints on the error signal reaching the retina, rather than the feedforward bottlenecks themselves.
-To test this hypothesis, we trained a model similar to that of Lindsey et.
-al  [29] but with full-width layers throughout.
-Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with RAF (Figure 4c), thereby isolating the impact of error dimensionality on neural representations.
+We hypothesize that these effects are due to constraints on the error signal reaching the retina rather than the feedforward bottlenecks themselves.
+To test this hypothesis, we trained a model similar to that of Lindsey et al.
+[29], but with full-width layers throughout.
+Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with LDFA (Figure 4c), thereby isolating the impact of error dimensionality on neural representations.
 We extracted the receptive fields of neurons in the retinal layer using visualization techniques [30] (Figure 4d).
 Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer.
-To further validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dRAF to train the retinal layer without restricting the backward pathway.
+To validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dLDFA to train the retinal layer without restricting the backward pathway.
 In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix).
@@
 Appendix A Full linear theory
-in this section, we provide the analysis of the linear networks for both Feedback Alignment and RAF
+in this section, we provide the analysis of the linear networks for both Feedback Alignment and LDFA
 Detailed analysis of Feedback Alignment learning dynamics
@@
 $\tau\frac{d\bar{W}_{2}}{dt}$ $=\left(S-\bar{W}_{2}\bar{W}_{1}\right)\bar{W}_{1}^{\top}.$
-Full derivation of Restricted Adaptive Feedback (RAF) learning dynamics
+Full derivation of low-dimensional feedback alignment (LDFA) learning dynamics
 We consider an alternative algorithm where the matrix $B$ is replaced by $B=QP$, with $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$.
@@
 Our implementation and optimization were conducted entirely using PyTorch.
-For the RAF layers, we initialized all weights using Kaiming uniform initialization.
+For the LDFA layers, we initialized all weights using Kaiming uniform initialization.
 We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule.
@@
 $\Delta W_{l}\leftarrow\boldsymbol{\delta}_{l+1}^{\top}\boldsymbol{h}_{l}$;
-Algorithm 1 Modified Backward Pass for RAF Layer
-The derivation of the dRAF (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
+Algorithm 1 Modified Backward Pass for LDFA Layer
+The derivation of the dLDFA (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
 Appendix C Numerical experiments
@@
 In Figure [3.a], we present the results of training a network with four hidden layers, each containing 512 neurons, and an output layer with 10 neurons on the CIFAR-10 dataset.
-The network was trained using RAF, as described earlier, without rank constraints, except for one layer at a time.
+The network was trained using LDFA, as described earlier, without rank constraints, except for one layer at a time.
 For comparison, we also trained the network using standard backpropagation as a baseline.
@@
 Each training run was repeated 5 times.
-dRAF
-For the results shown in Figure [3.d], we trained the same model using dRAF.
+Direct LDFA
+For the results shown in Figure [3.d], we trained the same model using dLDFA.
 In one experiment, we propagated the error signal from the last layer to all preceding layers.
@@
 Convolutional Neural Networks
-To extend our RAF algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal.
+To extend our LDFA algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal.
 This effectively constrains the dimensionality of the error signal across the spatial dimensions.
@@
 A bottleneck was introduced between the retinal and VVS layers to constrain the information flow.
-Training was performed using the Restricted Adaptive Feedback (RAF) method, where rank constraints were applied to feedback sent to the retina with ranks $r=2,4,32$.
+Training was performed using the low-dimensional feedback alignment (LDFA) method, where rank constraints were applied to feedback sent to the retina with ranks $r=2,4,32$.
 The network was trained with hyperparameters similar to those in Linsdey et.
@@
 Each model was trained for 120 epochs, and all experiments were repeated five times to ensure robustness.
-Additionally, we conducted experiments with dynamic Restricted Adaptive Feedback (dRAF), where feedback to the retina originated from higher visual layers while feedforward connections within the retina were constrained to four channels.
+Additionally, we conducted experiments with dynamic Restricted Adaptive Feedback (dLDFA), where feedback to the retina originated from higher visual layers while feedforward connections within the retina were constrained to four channels.
 Table C1: CNN Architecture

### v3

@@
 Significance statement
-Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron, a process with limited biological plausibility.
-Our work demonstrates that neural networks can be trained effectively using low-dimensional error signals, comparable to the inherent complexity of the task rather than the size of the network.
-By developing a novel, biologically plausible learning paradigm based on adaptive low-dimensional feedback, we show theoretically and empirically that minimal error dimensionality can achieve competitive performance even in complex neural architectures.
-These findings challenge existing assumptions about neural optimization and suggest fundamental constraints that may shape learning in both artificial and biological systems.
+Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron—a process lacking biological plausibility.
+Our work demonstrates that effective learning can instead occur with low-dimensional error signals, whose dimensionality scales with task complexity rather than network size.
+By developing a novel learning paradigm that employs local plasticity rules to adaptively align feedback signals with task-relevant errors, we show theoretically and empirically that even minimal error dimensionality achieves competitive performance in complex neural architectures.
+These findings challenge current assumptions about neural optimization and highlight fundamental constraints that can govern learning processes in both artificial and biological neural systems.
 1.   Introduction
@@
 Moreover, the brain likely operates with low-dimensional error signals, reflecting the inherently low-dimensional nature of tasks.
-This raises an essential question: Can large neural networks be trained efficiently with constrained, low-dimensional error signals?
+This raises an essential question: Can large neural networks be trained efficiently with partial information about the loss gradient carried by low-dimensional error signals?
 If error dimensionality is tied to task complexity rather than network size, it could revolutionize how we train large networks and improve our understanding of biological learning.
@@
 The strong coupling of feedforward and backward passes in backpropagation, where both processes use the same synapses, has limited investigations into alternative training methods.
-While approaches such as Feedback Alignment have been proposed [5], the necessity of full gradients for effective learning has not been questioned.
+Some approaches, such as Feedback Alignment [5], were proposed to address the problem of weight transposition in backporpagation, the necessity of full gradients for effective learning has not been questioned.
 In this study, we propose that high-dimensional error propagation is not strictly necessary for training deep, overparameterized networks.
@@
 The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$.
-(b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=64$ and $r=8$, respectively.
+(b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=m=64$ and $r=8<m$, respectively.
 The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$.
@@
 Specifically, the Johnson–Lindenstrauss lemma  [17] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal would be maintained.
-Nevertheless, our analysis shows that the solutions weight are not guaranteed to converge to the correct solution.
+Nevertheless, our analysis shows that the solutions weights are not guaranteed to converge to the correct solution.
 Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail.
+Furthermore, FA fails when $k<m$ even if the rank of $B$ is not restricted.
 The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.
@@
 Specifically, lower-dimensional error signals promote higher symmetries in emergent receptive fields.
-Our findings underscore the importance of considering the dimensionality and pathways of error signals when studying neural computations in the brain.
+Importantly, the architecture studied by Lindsey et al.
+is not a faithful model of the early visual system.
+In particular, there is no evidence that error signals propagate from the cortex or thalamus back to the retina.
+Instead, it is believed that the receptive fields in the retina emerge through unsupervised learning [31].
+Nevertheless, our results provide an explanation for the receptive fields observed by Lindsey et al.
+Furthermore, it may lead to an alternative explanation of learning in the visual pathway that uses simplified signaling (e.g., neuromodulators).
+Most importantly, our findings show a relation between tuning curves and error pathways.
 6.   Discussion

### v4

@@
-Training Large Neural Networks With Low-Dimensional Error Feedback
+Efficiently Training Large Neural Networks With Low-Dimensional Error Feedback
 Abstract
-Training deep neural networks typically relies on backpropagating high-dimensional error signals—a computationally intensive process with little evidence supporting its implementation in the brain.
-However, since most tasks involve low-dimensional outputs, we propose that low-dimensional error signals may suffice for effective learning.
-To test this hypothesis, we introduce a novel local learning rule based on Feedback Alignment that leverages indirect, low-dimensional error feedback to train large networks.
-Our method decouples the backward pass from the forward pass, enabling precise control over error signal dimensionality while maintaining high-dimensional representations.
-We begin with a detailed theoretical derivation for linear networks, which forms the foundation of our learning framework, and extend our approach to nonlinear, convolutional, and transformer architectures.
-Remarkably, we demonstrate that even minimal error dimensionality—on the order of the task dimensionality—can achieve performance matching that of traditional backpropagation.
-Furthermore, our rule enables efficient training of convolutional networks, which have previously been resistant to Feedback Alignment methods, with minimal error.
-This breakthrough not only paves the way toward more biologically accurate models of learning but also challenges the conventional reliance on high-dimensional gradient signals in neural network training.
-Our findings suggest that low-dimensional error signals can be as effective as high-dimensional ones, prompting a reevaluation of gradient-based learning in high-dimensional systems.
-Ultimately, our work offers a fresh perspective on neural network optimization and contributes to understanding learning mechanisms in both artificial and biological systems.
+Training large neural networks relies on propagating detailed error signals backward through every layer, a procedure that is computationally expensive and difficult to reconcile with biological teaching pathways.
+In the brain, instructive signals are typically sparse, indirect, and compressed, raising the question of whether a low-dimensional teaching signal can suffice for effective credit assignment without transmitting a full layerwise gradient.
+Building on Feedback Alignment, we introduce a low-dimensional feedback framework that decouples the backward and forward passes and explicitly controls the dimensionality of the error signal delivered to each layer while preserving high-dimensional feedforward representations.
+We develop a theory for the nonlinear learning dynamics induced by gradient descent in linear networks.
+Fixed, task-agnostic low-dimensional feedback can misassign credit and lead to spurious solutions, whereas allowing the feedback subspace to adapt toward task-relevant error directions recovers backpropagation-like learning dynamics.
+Guided by these predictions, we train convolutional networks and vision transformers to backpropagation-level performance with modest feedback dimensionality that scales with task dimensionality rather than network width, while reducing total training compute by over 20% without loss of accuracy.
+We further derive synapse-local rules that learn the relevant error subspace from pre- and postsynaptic signals, providing a biologically grounded mechanism for compressed teaching.
+Finally, feedback dimensionality reshapes receptive fields, revealing the teaching pathway as an inductive bias in representation learning.
 Significance statement
-Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron—a process lacking biological plausibility.
-Our work demonstrates that effective learning can instead occur with low-dimensional error signals, whose dimensionality scales with task complexity rather than network size.
-By developing a novel learning paradigm that employs local plasticity rules to adaptively align feedback signals with task-relevant errors, we show theoretically and empirically that even minimal error dimensionality achieves competitive performance in complex neural architectures.
-These findings challenge current assumptions about neural optimization and highlight fundamental constraints that can govern learning processes in both artificial and biological neural systems.
+Training modern neural networks often depends on sending detailed error information backward through every layer, a process that is costly in time, energy, and memory.
+In contrast, the brain appears to rely on simpler teaching signals.
+We show that accurate learning can be driven by a compact feedback signal whose size is set by task demands rather than by the number of neurons or synapses.
+Building on this principle, we develop a training framework that matches the accuracy of the standard training method while being more efficient.
+We also describe local learning rules for constrained teaching pathways and show that these constraints shape features learned in early layers, offering a new route for modeling learning in brain circuits.
 1.  Introduction
-The brain, despite its vast complexity, operates within the constraints of a low-dimensional world.
-Every movement—lifting a hand, turning a head—follows finite joint constraints.
-Navigation unfolds in modest spatial dimensions, and object recognition is drawn from a limited set of familiar categories.
-Even when possibilities seem vast, they are dwarfed by the sheer number of neurons and synapses in any brain region.
-Artificial neural networks reflect this paradox: the number of neurons in each layer often far exceeds the task-relevant variables.
-This contrast between task and internal representation dimensionality challenges us to rethink how we train such large networks and explore new learning paradigms.
-Modern deep learning methods rely on high-dimensional gradients that propagate backward through layers, ensuring a precise credit assignment to each parameter [1, 2].
-However, this process is computationally intensive and biologically implausible, given the lack of direct evidence for backpropagation in the brain [3, 4].
-If the loss gradient does not propagate back through the network, it must be obtained through alternative routes [5].
-In this case, there is no reason to assume that each neuron receives detailed information about the local gradient.
-Moreover, the brain likely operates with low-dimensional error signals, reflecting the inherently low-dimensional nature of tasks.
-This raises an essential question: Can large neural networks be trained efficiently with partial information about the loss gradient carried by low-dimensional error signals?
-If error dimensionality is tied to task complexity rather than network size, it could revolutionize how we train large networks and improve our understanding of biological learning.
-Despite the importance of this question, the effect of low-dimensional error feedback on learning and neural representations was never properly explored.
-The strong coupling of feedforward and backward passes in backpropagation, where both processes use the same synapses, has limited investigations into alternative training methods.
-Some approaches, such as Feedback Alignment [5], were proposed to address the problem of weight transposition in backporpagation, the necessity of full gradients for effective learning has not been questioned.
-In this study, we propose that high-dimensional error propagation is not strictly necessary for training deep, overparameterized networks.
-We introduce a novel learning rule based on Feedback Alignment that decouples forward and backward passes, enabling precise control over the dimensionality of error signals.
-Our method exploits the low-dimensional nature of real-world tasks to train large networks efficiently without compromising their representational capacity, achieving performance on par with backpropagation.
-A theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance.
-Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is feasible and universal.
-Finally, we explore how error dimensionality influences neural representations by examining its effects on receptive fields in an early visual system model, underscoring the broader implications of our findings.
+Training large neural networks depends on assigning credit across many layers and millions of parameters.
+The standard approach, gradient descent with error backpropagation, computes reverse-mode derivatives by transporting high-dimensional error information through every layer.
+This backward pass is not only a conceptual constraint but also a practical bottleneck because it requires manipulating and storing large intermediate signals whose size scales with hidden-layer width, leading to substantial memory, bandwidth, and compute costs during training [5, 11, 55].
+While large forward representations are needed for expressivity, it is not obvious that the feedback pathway must also scale with model size.
+This raises a basic efficiency question.
+Many tasks have low-dimensional outputs, so the loss gradient at the output layer is naturally low-dimensional, with an effective dimensionality tied to task complexity rather than to hidden-layer width.
+The open issue is whether deep networks can be trained by transmitting only a task-relevant, low-dimensional teaching signal to each layer, thereby reducing the overhead of the backward pass without sacrificing the benefits of high-dimensional feedforward representations.
+This efficiency-motivated perspective is also consistent with biological circuitry.
+Biological circuits do not exhibit the synapse-wise, bidirectional connectivity and precise weight symmetry required by gradient backpropagation.
+Instead, experimentally observed instructive channels are typically indirect, not originating from the direct downstream neural population, and are often sparse and long-range, which implies a restricted teaching signal [42].
+Notable examples include midbrain dopaminergic neurons [62, 45, 6], the olivocerebellar system [28, 47], cholinergic neuromodulation [33, 21, 36], and the noradrenergic system in the Locus Coeruleus [2, 7, 59, 67, 10].
+Combined with empirical evidence that task-relevant population activity is often confined to low-dimensional manifolds [12, 16, 57, 22], these observations suggest that the brain’s teaching pathway is more plausibly indirect and compressed (low-dimensional) than a neuron-by-neuron gradient trace.
+A parallel question arises in artificial learning.
+Existing biologically motivated alternatives to backpropagation, such as Feedback Alignment (FA) [38], decouple feedforward and feedback weights but still propagate full-dimensional error signals and often fall short of backpropagation in deep or complex architectures [4].
+As a result, the role of error dimensionality as an explicit constraint on deep credit assignment has been difficult to isolate.
+Here we develop a theoretical and algorithmic framework for training large networks with constrained, low-dimensional error feedback.
+We treat the error pathway as a controllable resource and ask which error subspace must be transmitted for effective learning, and how such a subspace can be learned with local plasticity rules.
+Our approach imposes a rank-$r$ constraint on feedback and learns a task-relevant low-dimensional teaching channel, thereby decoupling the backward pathway from the feedforward computation while preserving high-dimensional forward representations.
+We refer to this framework as low-dimensional feedback alignment (LDFA).
+Our work makes four contributions:
+  • We show that accurate credit assignment does not require full-dimensional backward signals.
+In a solvable linear theory, fixed low-rank feedback makes the stationarity conditions underdetermined and introduces spurious fixed points, and we characterize how the rank required for backpropagation-like convergence depends on task dimensionality.
+  • We introduce a factorized low-rank feedback pathway together with learning rules that make the dimensionality of the layerwise teaching signal an explicit, tunable resource.
+  • We demonstrate in deep nonlinear networks, including convolutional architectures and vision transformers, that near-backpropagation accuracy is achievable with ranks on the order of task complexity and largely independent of network width, and we quantify the resulting reduction in backward-pass compute.
+  • We show that error dimensionality systematically shapes learned representations, including receptive-field structure in a simplified ventral-stream model.
+Together, these results challenge the assumption that scalable credit assignment must involve transporting a full high-dimensional gradient through depth.
+Instead, they support a view in which learning can be driven by a learned low-dimensional teaching channel that preserves task-relevant error directions, yielding a computationally efficient training regime and a concrete hypothesis for how anatomically restricted feedback pathways could still support deep learning.
 2.  Background and related work
-We consider a multilayered perceptron with $L$ layers, each layer $l$ computes its output as $\boldsymbol{h}_{l}=f(W_{l}\boldsymbol{h}_{l-1})$, where $W_{l}$ is the weight matrix and $f$ is an element-wise activation function.
-The input to the network is $\boldsymbol{h}_{0}=\boldsymbol{x}$, and the final network output is $\hat{\boldsymbol{y}}=f_{L}(W_{L}\boldsymbol{h}_{L-1})$, which approximates the target $\boldsymbol{y}$.
-The task dimensionality, denoted $d$, is at most the number of components in $\boldsymbol{y}$ and $\hat{\boldsymbol{y}}$.
-Training the network involves minimizing a loss function $\mathcal{L}(\boldsymbol{y},\hat{\boldsymbol{y}})$ by adjusting the weights $\{W_{l}\}$.
-The error signal at the output layer, $\delta_{L}=\frac{\partial\mathcal{L}}{\partial\hat{\boldsymbol{y}}}$, is a $d$-dimensional vector, typically much smaller than the number of neurons in the hidden layers.
+[…]
 Backpropagation (BP) [56] is the standard approach for training neural networks.
-It propagates the error backward through the network using $\delta_{l}=W_{l+1}^{T}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1})$, and updates the weights using $\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{T}$, where $\eta$ is the learning rate.
-However, this method requires an exact transposition of the forward weights, $W_{l+1}^{T}$, which is biologically implausible [6, 7].
+It propagates the error backward through the network using $\delta_{l}=W_{l+1}^{\top}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1})$, and updates the weights using $\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{\top}$, where $\eta$ is the learning rate.
+However, this method requires an exact transposition of the forward weights, $W_{l+1}^{\top}$, which is biologically implausible [24, 15].
 Moreover, BP tightly couples the error propagation with the forward pass, limiting the ability to explore how different properties of the error signal affect learning dynamics.
-Feedback Alignment (FA) [5] was proposed to address the biological limitations of BP by replacing $W_{l+1}^{T}$ with a fixed random matrix $B_{l}$. The error is computed as:
+Feedback Alignment (FA) [38] was proposed to address the biological limitations of BP by replacing $W_{l+1}^{\top}$ with a fixed random matrix $B_{l}$. The error is computed as:
 $\delta_{l}=B_{l}\delta_{l+1}\odot f^{\prime}(W_{l}\boldsymbol{h}_{l-1}),$ (1)
@@
 An extension of FA involves adapting $B_{l}$ by updating it alongside the forward weights $W_{l}$ to improve their alignment [34, 1]:
-$\Delta B_{l}=-\eta\boldsymbol{h}_{l-1}\delta_{l}^{T}-\lambda B_{l},\quad\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{T}-\lambda W_{l},$ (2)
+$\Delta B_{l}=-\eta\boldsymbol{h}_{l-1}\delta_{l}^{\top}-\lambda B_{l},\qquad\Delta W_{l}=-\eta\delta_{l}\boldsymbol{h}_{l-1}^{\top}-\lambda W_{l},$ (2)
 where $\lambda$ is a regularization parameter.
-Although this adaptive approach improves performance by better aligning forward and backward weights, it still requires high-dimensional error signals and struggles to match BP performance in complex architectures like CNNs.
-Furthermore, in Section 3, we show that this approach fails when the matrix $B$ is low-rank and the dimensionality of the error is constrained.
+Although this adaptive approach improves performance by better aligning forward and backward weights, it still requires high-dimensional error signals and and is less efficient than backpropagation.
+Furthermore, in section˜3, we show that this approach fails when the matrix $B$ is low-rank and the dimensionality of the error is constrained.
 Other studies have explored the use of fixed sparse feedback matrices to reduce the dimensionality of error propagation [14].
@@
 However, no previous work has considered training with a constrained error pathway, and the effects of error dimensionality and training have not been systematically studied.
-In this work, our objective is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of large neural networks.
-To this end, we introduce a novel learning scheme that enables deep layers to be learned using low-rank feedback matrices, which allow flexible control over the dimensionality of the loss gradients in each layer (Figure 1).
-Our main contributions are:
-  1. 1.
-    We present a novel local learning rule for low-rank feedback weights, which matches BP performance while requiring minimal error signals.
-We provide a detailed derivation of the learning dynamics in a simple linear case, establishing the foundation for why and how the feedback weights should be learned.
-  2. 2.
-    We demonstrate that large and deep nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals and highlight that the feedback dimensionality should scale with the task rather than the size of the network.
-  3. 3.
-    We successfully train complex yet highly useful architectures, such as convolutional networks and visual transformers, using low-dimensional feedback.
-  4. 4.
-    We study a simplified model of the ventral visual system and show that the emerging receptive fields depend on the error-feedback dimensionality.
-These findings offer new insights into the relationship between learning mechanisms and biological neural representations.
-In the final section, we discuss the broader implications of our results for both neuroscience and machine learning.
+In this work, we ask how limiting the dimensionality of the propagated error constrains credit assignment in large neural networks.
+We treat the error pathway as a controllable resource and study how its dimensionality shapes learning dynamics, computational cost, and the representations that emerge during training.
+To do so, we introduce a learning scheme in which each layer receives error through a low-rank feedback map, enabling explicit control over the dimensionality of the layerwise teaching signal (fig.˜1).
+This focus differs from prior observations that learning dynamics or weight updates often exhibit low-dimensional structure: those results describe emergent compression but do not impose a bottleneck on the communicated error itself.
+The remainder of the paper develops a solvable linear theory of low-rank error feedback (section˜3), tests the predicted scaling in deep vision models (section˜4), derives synapse-local rules for learning the feedback pathway (section˜5), and then analyzes efficiency and representational consequences before closing with broader implications in the Discussion.
 Figure 1: Illustration of different approaches for propagating error to hidden layers.
-From left to right: Backpropagation (BP) propagates error using the exact transpose of the forward weights.
-Feedback Alignment (FA) replaces the transposed weights with fixed random feedback matrices, which gradually align with the forward weights during training.
-Our work framework provides low-dimensional feedback (LDF) using low-rank matrices, allowing full control of the dimnsionality of the gradients in each layer.
-Finally, direct low-dimensional feedback (dLDF) extends our LDF architecture by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
+From left to right: Backpropagation (BP) propagates error using the exact transposes of the forward weights.
+Feedback alignment (FA) replaces $W_{l}^{\top}$ with fixed random feedback matrices and relies on the forward weights aligning with this backward pathway during training.
+Low-dimensional feedback alignment (LDFA) constrains each feedback map to be low rank (e.g., $B_{l}=Q_{l}P_{l}$ with rank $r$), so each layer receives an $r$-dimensional teaching signal whose subspace is learned.
+Direct LDFA (dLDFA) extends LDFA by routing low-dimensional error projections to non-adjacent layers, allowing early layers to receive teaching signals directly from the output (or other higher) layers.
 3.  Low-dimensional error feedback in linear networks
@@
 Here, $\boldsymbol{x}\in\mathbb{R}^{n}$ represents the $n$-dimensional input, and $\boldsymbol{y}\in\mathbb{R}^{m}$ represents the target.
-The matrix $A$ is a rank-$d$ matrix defined as $A=\sum_{j=1}^{d}\boldsymbol{u}_{j}\boldsymbol{v}_{j}^{T}$, where $\boldsymbol{u}_{j}\in\mathbb{R}^{n}$ and $\boldsymbol{v}_{j}\in\mathbb{R}^{m}$ are random Gaussian vectors, and we assume $d\ll n$.
+The matrix $A$ is a rank-$d$ matrix defined as $A=\sum_{j=1}^{d}\boldsymbol{u}_{j}\boldsymbol{v}_{j}^{\top}$, where $\boldsymbol{u}_{j}\in\mathbb{R}^{n}$ and $\boldsymbol{v}_{j}\in\mathbb{R}^{m}$ are random Gaussian vectors, and we assume $d\ll n$.
 Our data set consists of $p$ training samples $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}_{\mu=1}^{p}$, with each input vector $\boldsymbol{x}^{\mu}$ being i.i.d.
@@
 We apply Feedback Alignment (FA) to update $W_{1}$, which does not have direct access to the loss gradient.
-Instead of backpropagating the error through $W_{2}^{T}$, we use a fixed low-rank feedback matrix $B$.
+Instead of backpropagating the error through $W_{2}^{\top}$, we use a fixed low-rank feedback matrix $B$.
 This provides an alternative pathway for propagating the error signal to $W_{1}$.
 For a given data point $\{\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu}\}$, the weight updates, derived from the FA framework, are given by:
-$\Delta W_{1}^{\mu}$ $=\eta B(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
-$\Delta W_{2}^{\mu}$ $=\eta(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T}W_{1}^{T}.$
+$\Delta W_{1}^{\mu}$ $=\eta B(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu\top},\text{ and }$ (4)
+$\Delta W_{2}^{\mu}$ $=\eta(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu\top}W_{1}^{\top}.$
 Here, $\eta$ represents the learning rate, and the update for $W_{1}$ is computed using the indirect error feedback provided by $B$, while $W_{2}$ receives the full error signal directly from the output.
@@
 [60] to incorporate indirect feedback with constrained dimensionality.
-We begin by characterizing the task across the $p$ data points through the input-output covariance matrix, $\Sigma_{io}=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}(\boldsymbol{x}^{\mu})^{T}$, which captures the correlation between input vectors $\boldsymbol{x}$ and output vectors $\boldsymbol{y}$.
-Performing Singular Value Decomposition (SVD), we obtain $\Sigma_{io}=USV^{T}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ contain the left and right singular vectors, respectively, and $S\in\mathbb{R}^{m\times n}$ is a rectangular diagonal matrix of singular values.
+We begin by characterizing the task across the $p$ data points through the input-output covariance matrix, $\Sigma_{io}=\frac{1}{p}\sum_{\mu=1}^{p}\boldsymbol{y}^{\mu}(\boldsymbol{x}^{\mu})^{\top}$, which captures the correlation between input vectors $\boldsymbol{x}$ and output vectors $\boldsymbol{y}$.
+Performing Singular Value Decomposition (SVD), we obtain $\Sigma_{io}=USV^{\top}$, where $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ contain the left and right singular vectors, respectively, and $S\in\mathbb{R}^{m\times n}$ is a rectangular diagonal matrix of singular values.
 For sufficiently large $p$, the first $d$ singular values in $S$ correspond to the prominent directions in the data (i.e., the singular values of $A$), while the remaining singular values are $O(1/\sqrt{p})$ and dominated by noise.
@@
 This transformation simplifies the analysis by aligning the network’s weight dynamics with the key data directions:
-$W_{1}=\bar{W}_{1}V^{T},\quad W_{2}=U\bar{W}_{2},\quad B=\bar{B}U^{T},$
+$W_{1}=\bar{W}_{1}V^{\top},\qquad W_{2}=U\bar{W}_{2},\qquad B=\bar{B}U^{\top},$
 where $\bar{W}_{1}$, $\bar{W}_{2}$, and $\bar{B}$ represent the transformed weight matrices.
@@
 Assuming a small learning rate $\eta\ll 1$ with full-batch updates, we express the weight updates in continuous time:
-$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
+$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),\qquad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{\top},$ (5)
 where $\eta=dt/\tau$.
@@
 By analyzing these equations, we can identify fixed points and evaluate their stability, providing insight into how the network converges and learns under constrained feedback.
-Figure 2, shows how the singular vectors of $W_{2}W_{1}$ align to the corresponding singular vectors of $\Sigma_{io}$
 Figure 2: Learning dynamics and component alignment in linear networks.
@@
 The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$.
-(b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=m=64$ and $r=8<m$, respectively.
-The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$.
-(d, e) Same as (b, c) but training both and $Q$ and $P$ using Eqs (7) and (8) .
-(f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully.
-(g) Training $P$ to align with the error ensures full recovery of the top $r$ singular components, improving convergence and alignment.
-Stationary solutions for the training.
-Training halts when the right-hand side of the weight update equations (5) vanishes, indicating that the dynamics have reached a stable fixed point.
-At this fixed point, the update equation for $\bar{W}_{1}$ leads to the condition:
-$\bar{B}(S-\bar{W}_{2}\bar{W}_{1})=0\implies S_{jj}\boldsymbol{B}_{:,j}=\sum_{i=1}^{m}\boldsymbol{B}_{:,i}(\bar{W}_{2}\bar{W}_{1})_{ij}\quad\forall j,$ (6)
-where $\boldsymbol{B}_{:,j}\in\mathbb{R}^{k}$ is the $j$-th column of $\bar{B}$, and $S_{jj}$ is the $j$-th singular value of $\Sigma_{io}$.
-This equation indicates that, at the stationary point, the weight products $\bar{W}_{2}\bar{W}_{1}$ must align with the singular modes of the data.
-However, since $B$ is of rank $r$, the feedback matrix $\bar{B}$ can only span at most $r$ independent directions.
-If $r=m$, the system has enough feedback dimensionality to align perfectly with the singular values in $S$, recovering the full structure of $\Sigma_{io}$ as demonstrated in Saxe et.
-al.
-[16].
-In this case, the training successfully converges to a unique solution where $\bar{W}_{2}\bar{W}_{1}=S$ (Figure 2b).
-Crucially, when $r<m$, the feedback matrix $\bar{B}$ lacks the sufficient rank to fully capture the $m$ independent directions in $S$.
-As a result, eq.
-(6) becomes under-determined, leading to potentially infinite solutions.
-This means that the trained weights $\bar{W}_{2}\bar{W}_{1}$ may not align with the true data structure encoded in $\Sigma_{io}$ (Figure 2c)
-The implications of the uderdetrimed fixed point solutions are surprising.
-Naively, one would expect that if the bottleneck is not too narrow, the projections of the error would maintain the necessary structure to guide the learning.
-Specifically, the Johnson–Lindenstrauss lemma  [17] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal would be maintained.
-Nevertheless, our analysis shows that the solutions weights are not guaranteed to converge to the correct solution.
+(b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), updating $Q$ with $P$ fixed [1], with $r=m=64$ and $r=8<m$, respectively.
+The $y$-axis reports the normalized mode overlap $\Lambda_{i}=u_{i}^{\top}W_{2}W_{1}v_{i}/S_{ii}$ for each singular component ii.
+(d, e) Same as (b, c) but training both $Q$ and $P$ using Eqs.
+(9).
+(f) With $P$ fixed, overlaps increase on average (bold), but the leading singular components do not reliably reach $\Lambda_{i}\approx 1$ at low rank.
+(g) Training $P$ aligns the feedback subspace with the evolving error, yielding near-complete recovery of the top-$r$ components and improved convergence.
+Stationary solutions for training.
+Training halts when the right-hand side of the weight dynamics in (5) vanishes, that is, when the system reaches a stationary point. At such a point, the update equation for $\bar{W}_{1}$ yields
+$\bar{B}\,(S-\bar{W}_{2}\bar{W}_{1})=0\iff S_{jj}\,\boldsymbol{B}_{:,j}=\sum_{i=1}^{m}\boldsymbol{B}_{:,i}\,(\bar{W}_{2}\bar{W}_{1})_{ij}\qquad\forall j,$ (6)
+where $\boldsymbol{B}_{:,j}\in\mathbb{R}^{k}$ is the $j$-th column of $\bar{B}$ and $S_{jj}$ is the $j$-th singular value of $\Sigma_{io}$.
+Equation (6) constrains the mismatch between the target map $S$ and the learned input–output map $\bar{W}_{2}\bar{W}_{1}$ only through the subspace spanned by the columns of $\bar{B}$.
+A convenient way to quantify how well the learned map captures each singular mode of the data is to project $\bar{W}_{2}\bar{W}_{1}$ into the singular-vector basis of $\Sigma_{io}=USV^{\top}$.
+We therefore define the (diagonal) mode overlap
+$\Lambda_{i}=\frac{u_{i}^{\top}(\bar{W}_{2}\bar{W}_{1})v_{i}}{S_{ii}}=\frac{(U^{\top}\bar{W}_{2}\bar{W}_{1}V)_{ii}}{S_{ii}},$ (7)
+where $u_{i}$ and $v_{i}$ are the $i$-th columns of $U$ and $V$.
+The overlap $\Lambda_{i}$ is a normalized measure of how strongly the learned input–output map expresses singular mode $i$ of $\Sigma_{io}$, with $\Lambda_{i}=1$ corresponding to perfect recovery of that mode.
+When $\mathrm{rank}(\bar{B})=r=m$, the feedback pathway constrains all $m$ directions, and the stationary conditions select the backpropagation solution $\bar{W}_{2}\bar{W}_{1}=S$, as in the exact linear-network dynamics of Saxe et al.
+[60] (Fig.
+2b).
+Crucially, when $r<m$, the constraint (6) is generically underdetermined because $\bar{B}$ can span at most $r$ independent directions.
+As a result, there can be infinitely many stationary solutions, and $\bar{W}_{2}\bar{W}_{1}$ need not match the full singular structure of $\Sigma_{io}$, so that some mode overlaps $\Lambda_{i}$ can deviate substantially from $1$ (Fig.
+2c).
+The implications of underdetermined fixed-point solutions are surprising.
+Naively, one would expect that if the bottleneck in the feedback is not too narrow, the projections of the error would maintain the necessary structure to guide the learning.
+Specifically, the Johnson–Lindenstrauss lemma  [31] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal is maintained.
+Nevertheless, our analysis shows that the solution weights are not guaranteed to converge to the correct solution.
 Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail.
-Furthermore, FA fails when $k<m$ even if the rank of $B$ is not restricted.
+Furthermore, FA fails when there is a bottleneck, i.e., $k<m,n$, even if the rank of $B$ is not restricted.
 The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.
 3.2 Training the feedback weights
-From the previous analysis, it is apparent that the feedback matrix $B$ must be trained for the learning dynamics to converge to the correct solution of Eq.
-(6).
-When $B$ is low-rank and fixed, the network lacks sufficient capacity to transmit the full error information, which can impede learning.
-One approach is to adopt a learning rule inspired by Kolen-Pollack, as in eq. (2). When $B=QP$, the updates are given by
-$\Delta W_{2}^{\mu}=\eta\boldsymbol{\delta}^{\mu}\boldsymbol{x}^{\mu T}W_{1}^{T}-\lambda W_{2}\quad\text{ and }\quad\Delta Q^{\mu}=\eta W_{1}\boldsymbol{x}^{\mu}\boldsymbol{\delta}^{\mu T}P^{T}-\lambda Q,$ (7)
-where $\boldsymbol{\delta}^{\mu}=\frac{\partial\mathcal{L}(\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu})}{\partial\boldsymbol{y}}=\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}$ is the error gradient for the $\mu$-th data point, $\eta$ is the learning rate, and $\lambda$ is the regularization parameter.
-Importantly, in this framework only the column space of $B$, or in our case $Q$ are updated.
-However, adopting this learning framework is insufficient when $P$ remains fixed and randomly initialized (Figure 2c).
-Since $P$ can project onto at most $r$ unique directions of the error, it may not align with the relevant error subspace.
-In this case, the network may converge to an incorrect solution, as indicated by the non-uniqueness of solutions to (6) when $m>r$.
-Ideally, we want $P$ to be an orthogonal matrix whose $r$ columns span the top $r$ principal directions of the output-output correlation matrix $\Sigma^{yy}$.
-This alignment ensures that the most significant components of the error are propagated back through the network.
-In cases where the output correlations are unknown, we can update $P$ using a modified Oja learning rule [18]:
-$\Delta P^{\mu}=\eta P\boldsymbol{y}^{\mu}\boldsymbol{y}^{\mu T}(I-P^{T}P)-\lambda P,$ (8)
-where $I$ is the identity matrix. This rule adjusts $P$ incrementally so that its columns converge to the top $r$ principal components of the outputs $\{\boldsymbol{y}^{\mu}\}$.
-By training both $Q$ and $P$, we allow the feedback matrix $B=QP$ to adaptively align with the relevant error directions, enabling the network to learn the correct mappings even under constrained feedback dimensionality.
-Repeating the linear analysis that led to (5), we extend the derivation to our case with adaptive feedback weights.
-We define the transformed feedback matrix as $P=\bar{P}U^{T}$.
-This transformation aligns the feedback matrix $P$ with the principal components of the data, simplifying the analysis.
-Taking the continuous-time limit (with $\eta\to 0$ and $\eta p=\tau$), we obtain a set of differential equations that describe the learning dynamics of the forward and backward weights:
-$\tau\frac{d\bar{W}_{1}}{dt}$ $=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),$ (9)
-$\tau\frac{d\bar{W}_{2}}{dt}$ $=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T}-\lambda\bar{W}_{2},$
-and
-$\tau\frac{dQ}{dt}$ $=\bar{W}_{1}(S-\bar{W}_{2}\bar{W}_{1})^{T}\bar{P}^{T}-\lambda Q,$ (10)
-$\tau\frac{d\bar{P}}{dt}$ $=\bar{P}SS^{T}(I-\bar{P}^{T}\bar{P})-\lambda\bar{P}.$
-Here, $\bar{B}=Q\bar{P}$ and $\lambda$ is the regularization parameter.
-Note that $Q$ is not affected by rotation, as it does not come in contact with either the input or the output.
-The full derivation can be found in the Appendix.
-Notably, the updates to the feedback weights are local and follow learning rules that were well-studied in theoretical neuroscience [18, 19, 20, 21].
-Local plasticity makes our framework an attractive alternative for backpropagation in brain circuit models.
-We refer to this learning framework as low-dimensional feedback alignment (LDFA).
-Figure 2 compares the learning dynamics of LDFA with those of BP and FA where only the $Q$ matrix is learned.
-The results demonstrate that LDFA effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality.
-Furthermore, LDFA will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
-3.3 Restricted adaptive feedback in deep architectures
-Our weight-update equations can be naturally extended to deeper networks.
-In the single hidden layer model above, the backward weights $P$ were updated using the true labels $\boldsymbol{y}^{\mu}$.
-However, in deeper models, hidden layers do not have ground-truth representations.
-Instead, each layer relies on the local error signal, which propagates through the network according to $\boldsymbol{\delta}_{l}^{\mu}=Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}$.
-This error signal $\boldsymbol{\delta}_{l}^{\mu}$ provides the necessary information for learning at the layer $l$.
-To update the feedback weights $P_{l}$ in the absence of ground-truth representations, we use local error signals $\boldsymbol{\delta}_{l+1}^{\mu}$.
-As in the single-layer model, we use Oja’s rule to adjust $P_{l}$ to span the principal components of the error at the next layer, $\boldsymbol{\delta}_{l+1}^{\mu}$, ensuring efficient error propagation.
-The complete update rules for layer $l$ are given by
-$\Delta W_{l}^{\mu}&=\eta Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{h}_{l}^{\mu T}-\lambda W_{l}\end{aligned}\qquad\text{and}\qquad\begin{aligned} \Delta Q_{l}^{\mu}&=\eta\boldsymbol{h}_{l}^{\mu}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}-\lambda Q_{l},\\ \Delta P_{l}^{\mu}&=\eta P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu T}(I-P_{l}^{T}P_{l})-\lambda P,\end{aligned}$ (11)
-where $\eta$ is the learning rate, $\lambda$ is the regularization parameter, $\boldsymbol{h}_{l}^{\mu}$ is the activation of layer $l$, and $\boldsymbol{\delta}_{l+1}^{\mu}$ is the error signal from the next layer.
-By updating $P_{l}$ using the error signals, we ensure that the feedback weights of each layer are adapted to capture the most relevant directions in the error space, facilitating effective learning throughout the network.
-Notably, while we use the same learning rate $\eta$ and weight decay $\lambda$ for all components $\{W_{l}\}$, $\{Q_{l}\}$, and $\{P_{l}\}$, it can potentially differ.
-In the interest of brevity, we omit simulations of deep linear networks, as the extension from the single-layer case is straightforward.
-Instead, we proceed directly to deep nonlinear networks, where the impact of constrained error feedback presents more complex and interesting dynamics.
-4.   Nonlinear networks and complex architectures
-Adapting our low-dimensional feedback alignment (LDFA) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable.
-The local update rules in eq.
-(11) remain the same; the primary difference lies in the introduction of nonlinear activation functions during the propagation of signals and errors.
-Specifically, the forward and backward passes are modified as follows:
-$h_{l}=f(W_{l}h_{l-1}),\quad\delta_{l}=Q_{l}P_{l}\delta_{l+1}\odot f^{\prime}(h_{l}),$ (12)
-where $f$ is the nonlinear activation function applied element-wise, and $f^{\prime}$ is its derivative.
-To test whether our learning rule extends effectively from linear to nonlinear models, we trained deep networks on the CIFAR-10 dataset.
-We used a simple nonlinear model with four fully connected layers of 512 ReLU neurons each.
-While not state-of-the-art, this model provides a suitable testbed for evaluating our theory’s applicability to nonlinear architectures and complex data.
-To isolate the impact of feedback dimensionality, we applied the LDFA rule in eq.
-(11), constraining the rank of feedback matrices in one specific layer at a time while leaving the others unrestricted.
-We then measured the network’s test accuracy and compared it to backpropagation (BP) as a baseline.
-Figure 3a shows the accuracy as a function of the feedback rank $r_{l}$, with each curve representing a different constrained layer.
-Consistent with our findings for the linear model, constraining the feedback rank to $r=d=10$, in any layer, match BP performance (Figure 3a), where $d$ is the number of classes.
-Interestingly, the shallower layers performed well even under tighter rank constraints ($r_{l}<d$), suggesting that the deeper layers compensate for the limited feedback in the earlier layers by effectively adjusting their weights.
-This compensatory effect is possible because no information is lost during the feedforward pass, unlike in a bottlenecked network.
-To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of LDFA-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b).
-The results confirm that LDFA-trained networks leverage their width to maintain high performance despite feedback constraints.
-Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using LDFA can still match BP performance, demonstrating LDFA’s robustness even with dimensional constraints across the entire network.
-Task dimensionality determines the minimal rank
-Our linear analysis suggests that the error signal dimensionality needed for effective learning is tied to the loss gradient dimensionality, which depends on the number of classes in the data.
-To test this, we trained networks on subsets of CIFAR-100 with 50, 75, and 100 classes, constraining the ranks of all feedback matrices to $r$ (Figure 3c).
-The results show that network performance matches BP when the feedback rank equals the number of classes ($r=d$).
-It indicates that the minimal rank required for effective learning aligns with the task’s complexity, defined by the number of output classes.
-Low-dimensional error signal from the top layer
-Our theory shows how to propagate error signals from deeper layers to shallower ones using restricted adaptive feedback.
-However, error projections can also bypass intermediate layers entirely, leading to different variants of direct low-dimensional feedbackalignment (dLDFA), analogous to Direct Feedback Alignment [22].
-For example, we can make direct connections from the output or the penultimate layer to earlier layers, training these connections using our algorithm (Figure 3d).
-As with LDFA, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$.
-This model is particularly important because it is more flexible and has greater potential to explain learning in the brain, where error signals may arrive from different pathways.
-Figure 3: Low-dimensional feedback alignment (LDFA) efficiently trains nonlinear networks on CIFAR-10.
-(a) Constraining feedback in any layer to $r=d=10$ matches full BP performance.
-Different curves ($B_{1}$, $B_{2}$, and $B_{3}$) correspond to restricting feedback at different layers.
-(b) Reducing feedback dimensionality has minimal impact on performance, whereas reducing network width significantly degrades accuracy, indicating that high-dimensional representations are still utilized.
-(c) Minimal error dimensionality is sufficient for learning, as shown by subsampling classes from CIFAR-100.
-All feedback matrices are constrained to the same rank $r$.
-(d) Direct LDFA: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance.
-(e) Performance of a 4-block VGG-like convolutional network trained with LDFA, constraining the layers with 512 channels.
-Inset: Training the same network with all layers constrained to a fraction of their size; the smallest layer has 64 channels.
-(f) A transformer-based architecture trained to classify images, demonstrating that LDFA generalizes to more advanced network architectures.
-See Supplemental Information for details on all implementations.
-Convolutional neural networks
-Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even with complex datasets, matching backpropagation (BP) performance.
-Here, we extend this investigation to convolutional architectures.
-Training convolutional networks with Feedback Alignment (FA) is notoriously difficult [8, 23].
-Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using LDFA.
-We aim to determine whether convolutional networks can also benefit from low-dimensional error feedback.
-We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see the appendix for details).
-Using LDFA, we decoupled error propagation from the feedforward pass in all layers.
-Initially, we constrained the feedback error only in the blocks containing 512 (Figure 3e).
-Consistent with our findings in fully connected networks, convolutional networks learn well with a feedback matrix with a rank similar to the number of classes, $d=10$.
-To test whether convolutional networks can train when constraining all feedback paths, we further restricted each block to have feedback matrices with ranks equal to $1/2$, $1/4$, or $1/8$ of the block width (Figure 3e inset).
-Our results indicate that reducing the error dimensionality has a minimal impact on performance, except in the most extreme case.
-Specifically, constraining the feedback rank to $1/8$ of the block width resulted in a noticeable drop in performance.
-This finding aligns with our previous results, as the layer with 64 channels received feedback with a rank of $r=8$.
-Overall, our results demonstrate that convolutional networks can be efficiently trained using a minimal error signal.
-Transformers
-Transformers have become a dominant architecture in natural language processing and, more recently, have achieved state-of-the-art performance in image classification tasks [26, 27].
-Unlike convolutional and standard feedforward architectures, transformers rely on a self-attention mechanism that computes interactions between tokens through the multiplication of key, query, and value matrices [28].
-These operations introduce complex, nonlinear dependencies that make it unclear whether low-dimensional error feedback, as proposed in our framework, can effectively support learning in transformer models.
-To investigate this question, we applied our approach to visual transformer architectures trained for image recognition on the CIFAR-10 dataset.
-Each weight matrix in the transformer architecture, including key, query, and value matrices, was trained using low-dimensional error feedback from the subsequent layer.
-Feedback matrices were trained using the LDFA framework (see the appendix for details).
-The network’s performance as a function of the feedback rank, which was kept identical for all matrices, is shown in Figure 3f.
-Our results indicate that, consistent with our findings in both nonlinear and convolutional networks, an error signal with dimensionality as low as the task dimensionality (i.e.
-the number of output classes) is sufficient to train transformer architectures effectively (Figure 3f).
-Despite the intricate nature of self-attention computations, training with low-dimensional feedback yielded performance comparable to conventional backpropagation on CIFAR-10 with rank $r=10$.
-Notably, training transformers with low-dimensional feedback using LDFA not only matched but, in some cases, consistently outperformed backpropagation by a small margin.
-This improvement was observed even when hyperparameters were independently optimized for both BP and LDFA.
-The improved generalization could be understood as a regularization effect achieved by constraining the feedback.
-However, since we observe it only in the transformer architecture, it is not a general property of a low-dimensional error signal, and we leave further investigation to a future study.
-Figure 4: Receptive fields (RFs) are shaped by feedback dimensionality rather than feedforward constraints.
-(a) Model of the early ventral visual stream.
-Top: Retinal-to-cortical connectivity is modeled as a bottleneck in the network; adapted from [29].
-Middle: A bottleneck is applied only to the feedback connections, restricting error dimensionality.
-Bottom: Using direct LDFA (dLDFA), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck.
-(b) Receptive fields (RFs) in a model trained with BP, corresponding to the setup in (a) top.
-Left: No bottleneck.
-Right: Bottleneck at the retinal output, which induces center-surround RFs.
-(c) Training the retinal output with LDFA.
-Restricting the rank of the feedback matrix produces center-surround RFs.
-(d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dLDFA.
-The presence of high-dimensional error signals results in oriented RFs.
+As shown above, when $\mathrm{rank}(B)=r<m$ the stationarity condition in Eq.
+(6) is underdetermined, and fixed low-rank feedback (e.g., random $B$) can drive the dynamics to incorrect fixed points.
+We therefore keep the rank constraint, but allow the feedback pathway to adapt by parameterizing $B=QP$ with $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$.
+The target feedback map is not known a priori, but backpropagation provides an ideal reference: exact credit assignment is obtained for $B=W_{2}^{\top}$.
+We thus learn the best rank-$r$ approximation of $W_{2}^{\top}$ by minimizing the factorization loss
+$\mathcal{L}_{B}=\frac{1}{2}\|QP-W_{2}^{\top}\|_{F}^{2}.$ (8)
+Notably, $\mathcal{L}_{B}$ involves only $W_{2}$ and the feedback parameters, so it is independent of the data and does not scale with batch size.
+Gradient descent on $\mathcal{L}_{B}$ yields
+$\Delta P$ $=Q^{\top}\bigl(W_{2}^{\top}-QP\bigr),$ (9)
+$\Delta Q$ $=\bigl(W_{2}^{\top}-QP\bigr)P^{\top}.$
+Thus, the feedback updates do not require access to activities or error signals; their only coupling to the training problem is through the evolving forward weights $W_{2}$.
+As in the analysis above, we rotate into the singular basis of $\Sigma_{io}$ by writing $P=\bar{P}U^{\top}$ and $W_{2}=U\bar{W}_{2}$, where $U$ contains the left singular vectors of $\Sigma_{io}$.
+Taking the continuous-time limit (with learning rate $\eta_{B}=dt/\tau_{B}$) gives the mode-wise dynamics
+$\tau_{B}\frac{d\bar{P}}{dt}$ $=Q^{\top}\bigl(\bar{W}_{2}^{\top}-Q\bar{P}\bigr),$ (10)
+$\tau_{B}\frac{dQ}{dt}$ $=\bigl(\bar{W}_{2}^{\top}-Q\bar{P}\bigr)\bar{P}^{\top}.$
+Together with the forward dynamics, Eqs.
+(10) define a closed system in which feedforward and feedback weights co-adapt.
+As shown in fig.˜2c,d, learning both $P$ and $Q$ eliminates the spurious low-rank fixed points and restores convergence to the correct solution under constrained feedback.
+4.  Low-dimensional error in deep vision models
+The linear theory above isolates the key obstruction induced by fixed low-rank feedback: when $\mathrm{rank}(B)=r<m$, the stationarity condition becomes underdetermined and learning can converge to spurious fixed points.
+Learning the feedback weights removes this degeneracy by making the backward mapping track the evolving feedforward transposes.
+Here we show that the same mechanism supports effective training in modern deep vision architectures, and we quantify the resulting scaling and computational consequences.
+This extension is conceptually direct.
+The feedback factors are updated only through the mismatch $|Q_{l}P_{l}-W_{l}^{\top}|_{F}^{2}$, which depends on the feedforward weights but not on activities, batch statistics, or the sampled errors used to train $W_{l}$.
+Consequently, inserting local nonlinearities in the forward pass does not alter the feedback-learning rule.
+We first state the multilayer nonlinear construction explicitly, and then apply it to convolutional and transformer-based vision models.
+4.1 Multilayer nonlinear perceptrons
+We consider an $L$-layer multilayer perceptron with activities $\boldsymbol{h}0=\boldsymbol{x}$ and
+$\boldsymbol{a}_{l+1}=W_{l}\boldsymbol{h}l,\hskip 18.49988pt\boldsymbol{h}_{l+1}=f(\boldsymbol{a}_{l+1}),\hskip 18.49988ptl=0,\ldots,L-1,$ (11)
+where $f$ is applied element-wise.
+Let $\boldsymbol{\delta}_{l}$ denote the backpropagated error at layer $l$, including the derivative of the nonlinearity.
+Under low-dimensional feedback, we replace the transpose $W_{l}^{\top}$ in the backward recursion with a learned rank-$r$ factorization $B_{l}=Q_{l}P_{l}$, giving
+$\boldsymbol{\delta}_{l}=(Q_{l}P_{l}\boldsymbol{\delta}_{l+1})\odot f^{\prime}(\boldsymbol{a}_{l}),\hskip 18.49988ptl=1,\ldots,L-1,$ (12)
+with the usual initialization at the output layer determined by the loss.
+Given these error signals, the feedforward weights follow the standard local outer-product update,
+$\Delta W_{l}^{\mu}=\eta\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{h}_{l}^{\mu\top}-\lambda W_{l},$ (13)
+where $\lambda$ is an $L_{2}$ regularization (weight decay) parameter.
+Finally, the feedback factors are trained to track the evolving transposes by gradient descent on $\frac{1}{2}|Q_{l}P_{l}-W_{l}^{\top}|_{F}^{2}$, yielding
+$\Delta P_{l}=$ $Q_{l}^{\top}\big(W_{l}^{\top}-Q_{l}P_{l}\big),\qquad\text{and}$ (14)
+$\Delta Q_{l}=$ $\big(W_{l}^{\top}-Q_{l}P_{l}\big)P_{l}^{\top}.$
+These updates depend only on $W_{l}$ and can therefore be computed without storing activations or using training data; in practice, they may also use a separate step size and can be applied on a slower update schedule than the forward weights.
+4.2 Deep convolutional networks
+Training convolutional networks with Feedback Alignment (FA) is known to be challenging [4, 35].
+Recent work improves FA by learning feedback weights [3], but our approach leverages a low-rank parameterization $B=QP$ that adapts both the output subspace that carries the propagated error (through $Q$) and the input subspace from which that error is assembled (through $P$).
+This factorized learning rule makes it possible to test whether deep convolutional networks can be trained with an error signal whose dimension is far smaller than the feedforward width, here measured by the number of channels per layer.
+We trained a VGG-like convolutional network [63] with four blocks and batch normalization on CIFAR-10 (see Appendix for details).
+In all layers, we replaced transpose feedback with learned low-rank backward matrices.
+Within each block, we fixed the feedback rank to be a fraction of the block width: $1/2$, $1/4$, or $1/8$ (fig.˜3a).
+Reducing the error dimensionality had little effect on test performance, except in the most extreme setting.
+When the rank was $1/8$ of the block width, performance dropped noticeably, consistent with the linear prediction that accuracy degrades when $r<d$: in our smallest block (64 channels), rank-$1/8$ feedback corresponds to $r=8$, below the CIFAR-10 task dimensionality $d=10$.
+Overall, these results show that deep convolutional networks can be trained effectively using a strongly compressed error signal.
+4.3 Transformer-based architectures
+Transformers achieve state-of-the-art performance in image classification [18, 32].
+Unlike convolutional and standard feedforward networks, they rely on self-attention, which mixes token representations through multiplicative interactions between key, query, and value projections [68].
+These interactions introduce strong cross-token dependencies, so it is not a priori clear that strongly compressed error signals provide sufficient credit assignment for learning in transformer models.
+To test this, we applied our low-rank feedback scheme to a vision transformer trained on CIFAR-10.
+All weight matrices in the architecture, including the key, query, and value projections, were trained using low-dimensional error feedback from the subsequent layer (see Appendix for details).
+For simplicity, we used the same feedback rank for all matrices, so the error dimensionality is controlled solely by the rank $r$ of each learned backward map.
+Performance as a function of feedback rank is shown in fig.˜3b.
+Despite the complexity of self-attention, transformers remained trainable with error dimensionality as low as the task dimensionality.
+In particular, rank $r=24$ feedback yielded performance on CIFAR-10 comparable to standard backpropagation (Figure 3b,c), consistent with our observations in linear and convolutional networks.
+4.4 Scaling and efficiency gains
+The results above show that low-dimensional feedback suffices to train deep nonlinear architectures, including convolutional networks and transformers, to near backpropagation-level accuracy.
+This has two implications: it decouples how performance scales with feedforward width from how the feedback pathway must scale, and it opens a regime where training can be more computationally efficient than standard backpropagation.
+Scaling with network width.
+Constraining the feedback to be low-rank changes the credit-assignment channel but leaves the feedforward model class unchanged.
+The representational capacity remains set by the width and depth of the forward network, not by the rank of the feedback matrices.
+Consistent with this, increasing the size of the feedforward network improves accuracy even when the feedback rank is held fixed (fig.˜3c top).
+At the same time, our linear analysis above indicate that the rank required to match backpropagation accuracy is controlled primarily by task dimensionality $d$ (e.g., performance degrades when $r<d$).
+Together, these observations imply that one can benefit from overparameterized feedforward representations while keeping the feedback dimensionality tied to task complexity rather than to network width.
+Computational efficiency.
+A direct benefit of low-rank feedback is a reduced cost of propagating errors through depth.
+For a layer with widths $n_{l}$ and $n_{l+1}$, backpropagation multiplies by $W_{l}^{\top}$ at cost $O(n_{l}n_{l+1})$ per example (or per token).
+With low-rank feedback $B_{l}=Q_{l}P_{l}$ of rank $r$, the same operation factors into $P_{l}\boldsymbol{\delta}{l+1}$ followed by $Q_{l}(\cdot)$, at cost $O(r(n_{l}+n_{l+1}))$.
+Accordingly, the FLOPs associated with backward error propagation decrease linearly with the effective error dimensionality $r$.
+Total training cost reflects three additional effects.
+First, the factors $Q_{l}$ and $P_{l}$ must be updated, but these updates depend only on $W_{l}$ and do not scale with batch size, so they can be amortized by updating the feedback factors at longer intervals.
+Second, extremely small ranks can slow convergence; in the regime where final performance matches backpropagation, learning curves are comparable (fig.˜3c top), whereas at very low ranks, additional epochs may be required.
+Third, the backward pass is only part of training, which also includes the full-rank forward pass and the weight updates.
+To capture these effects, we measure the total number of FLOPs required to reach $90\%$ of the final test accuracy, jointly accounting for backward-pass savings, feedback-factor updates, and any change in convergence time.
+This yields a U-shaped dependence on rank: at large $r$, per-iteration cost approaches backpropagation, whereas at very small $r$ slower convergence dominates.
+For a vision transformer on CIFAR-10, an intermediate rank achieves the best trade-off, reducing total compute (by about $20\%$ in our experiments) while maintaining backpropagation-level accuracy (fig.˜3c bottom).
+Figure 3: Low-dimensional feedback in nonlinear vision architectures.
+(a) Convolutional networks.
+A VGG-like CNN trained on CIFAR-10 with learned low-rank feedback achieves near-backpropagation (BP) test accuracy even when the backward/teaching channel is strongly compressed.
+Bars sphow final accuracy for BP and for low-dimensional feedback with feedback rank set to a fraction of the layer width (r=n/2,n/4,n/8r=n/2,\,n/4,\,n/8, where $n$ denotes the number of channels in the corresponding block).
+(b) Vision transformers (ViT).
+Top, schematic of a transformer block and multi-head self-attention with low-rank feedback channels; low rank feedback used to train all weights, including MLP and attention weights.
+Bottom, training curves (test accuracy versus training step) for BP (black) and low-dimensional feedback with a shared feedback rank $r$ across all linear maps (color-coded); inset shows the corresponding training loss.
+(c) Scaling and efficiency in ViT.
+Top, holding the feedback rank fixed ($r=36$) while reducing the embedding dimension decreases accuracy, indicating that performance continues to benefit from larger feedforward representations even under a fixed low-dimensional error channel.
+Bottom, ViT test accuracy (bars) and total compute required to reach $90\%$ of the final accuracy (line, TFLOPs) as a function of feedback rank, revealing an intermediate-rank regime that preserves BP-level accuracy while reducing end-to-end training compute (about $20\%$ in these experiments).
+5.  Local learning rules
+The factorized feedback-learning scheme above achieves backpropagation-level performance with an error channel of rank $r\ll n_{l}$, and yields tangible compute savings.
+However, at its core it remains a surrogate for weight transport since the update of $(Q_{l},P_{l})$ is driven by the corresponding forward matrix $W_{l}$ through the objective $\|Q_{l}P_{l}-W_{l}^{\top}\|_{F}^{2}$.
+Implementing such “learning through distance” would require synapses in the feedback pathway to access parameters in a separate, remote pathway, which is difficult to reconcile with known biological plasticity mechanisms.
+Moreover, explicitly approximating $W_{l}^{\top}$ presupposes a parallel forward route whose weights are being transposed; this assumption breaks down when error signals arrive through indirect pathways that do not mirror the feedforward circuit, as in direct feedback alignment [46].
+These limitations motivate a complementary perspective: rather than learning a proxy for $W_{l}^{\top}$, we seek a genuinely local mechanism that learns a useful low-dimensional error representation while keeping the remaining feedback circuitry fixed.
+Our linear analysis suggests that, under a rank constraint, the crucial degree of freedom is not an element-wise approximation of $W_{l}^{\top}$ but the error subspace that the feedback pathway can represent and transmit.
+Writing $B_{l}=Q_{l}P_{l}$, the matrix $P_{l}\in\mathbb{R}^{r\times m}$ projects output-space errors into an $r$-dimensional channel, so the columns of $P_{l}^{\top}$ span the subspace of output error directions that can be assembled upstream.
+When this subspace is poorly aligned with the task, the stationarity condition becomes underdetermined and training can converge to spurious fixed points.
+Conversely, when $P_{l}^{\top}$ spans task-relevant directions, the feedback pathway preserves informative components of the error while discarding directions that are redundant for learning.
+Ideally, $P_{l}$ has $r$ orthonormal rows ($P_{l}P_{l}^{\top}=I_{r}$), so that $P_{l}^{\top}P_{l}$ is the rank-$r$ projector onto the transmitted subspace in the downstream error space.
+Thus, $P_{l}$ learns the top-$r$ subspace of the incoming teaching signal to layer $l$ (i.e., the vector that $P_{l}$ multiplies).
+Concretely, we want the transmitted subspace to align with the top principal components of the layerwise error, given by the eigenvectors of $\frac{1}{p}\sum_{\mu}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu\top}$, where $\boldsymbol{\delta}_{l+1}^{\mu}$ denotes the locally available teaching/error signal layer $l+1$ for example $\mu$.
+Specifically, at the output layer, the error is given by $\boldsymbol{\delta}_{L}^{\mu}=\partial\mathcal{L}^{\mu}/\partial\hat{\boldsymbol{y}}^{\mu}$, while for hidden layers it is the error signal broadcast through the constrained feedback pathway.
+This choice allocates the limited rank of the feedback channel to the directions along which the network is currently making the biggest mistakes, and it uses a signal that is locally available wherever feedback arrives.
+To learn the principal components of the error using online updates, we use a multi-component Oja rule driven by the locally available teaching signal [48],
+$\Delta P_{l}^{\mu}=\eta\,P_{l}\,\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu\top}\,(I-P_{l}^{\top}P_{l})-\lambda P_{l},$ (15)
+where $I$ is the identity matrix.
+Eq.
+(15) is a matrix-valued Oja update for online principal-subspace learning.
+A Hebbian term driven by the teaching signal $\boldsymbol{\delta}_{l+1}^{\mu}$ increases the components of $P_{l}$ along directions of high error variance, while the remaining factors stabilize the dynamics and set the overall scale.
+In particular, the factor $(I-P_{l}^{\top}P_{l})$ promotes row-orthonormality of $P_{l}$, and the decay term $\lambda P_{l}$ controls the magnitude of the weights.
+Although the orthonormalizing factor is written compactly as a matrix product, its role is to express within-population competition and normalization in the $r$-dimensional error channel.
+In Oja’s original single-component rule, the stabilizing term can be interpreted as a postsynaptic-activity–dependent normalization of an otherwise Hebbian increment, yielding an online update that converges to the leading principal component [48].
+Extending PCA learning to multiple components necessarily requires a mechanism that prevents different components from collapsing onto the same direction, which is achieved by introducing competition among the units representing the learned subspace.
+Several biologically motivated constructions realize this competition using only local interactions within the postsynaptic population [58, 29, 30, 44, 27, 51].
+In particular, Sanger’s generalized Hebbian algorithm implements multi-component PCA by combining Hebbian plasticity with within-layer interactions that subtract already-learned projections, thereby enforcing orthogonality through local competition among units [58].
+More generally, similarity-matching formulations derive networks in which feedforward synapses learn Hebbianly while lateral synapses learn anti-Hebbianly, and the lateral dynamics locally decorrelate and normalize neural activities, effectively playing the role of the orthonormalization implicit in (15[51, 52].
+In this sense, (15) should be read as a compact description of a circuit-level computation in which synaptic updates depend on pre- and postsynaptic activity, while orthogonality and normalization are maintained by local competition within the error-channel population.
+Finally, Because $P_{l}$ adapts gradually, it tracks the dominant modes of the nonstationary teaching signal; the forward weights then reduce those modes, causing the subspace to rotate as learning progresses.
+Together with the layer-wise weight update (13), this yields a local learning scheme in which feedback synapses update from signals available at their pre- and postsynaptic terminals, and stabilization is mediated by within-layer normalization and competition rather than by nonlocal parameter access.
+We refer to this scheme as low-dimensional feedback alignment (LDFA).
+In what follows, we show that learning $P_{l}$ alone, with $Q_{l}$ fixed and random, is sufficient to reach backpropagation-level accuracy; as in Feedback Alignment [38], the forward weights can then align to the fixed mixing induced by $Q_{l}$.
+5.1 Linear analysis
+We analyze the coupled learning dynamics in the one-hidden-layer linear network of Section 3, using the same rotation into the singular basis of the input-output correlation $\Sigma_{io}=USV^{\top}$.
+Writing $W_{1}=\bar{W}_{1}V^{\top}$, $W_{2}=U\bar{W}_{2}$, and $P=\bar{P}U^{\top}$ yields an effective feedback $B=QP=Q\bar{P}U^{\top}$, with $Q$ fixed and $\bar{P}(t)$ learned.
+In the continuous-time limit, the forward dynamics in the rotated basis are exactly those of eq.˜5, with the sole modification that $\bar{B}(t)=Q\bar{P}(t)$ is now time-dependent.
+The new ingredient is the evolution of $\bar{P}$ under the error-driven Oja rule (15).
+For squared loss and whitened inputs, the output-layer error in the rotated coordinates satisfies $\bar{\boldsymbol{\delta}}^{\mu}(t)=\bar{E}(t)\,\bar{\boldsymbol{x}}^{\mu}$, where $\bar{\boldsymbol{x}}^{\mu}=V^{\top}\boldsymbol{x}^{\mu}$ and $\bar{E}(t)=S-\bar{W}_{2}(t)\bar{W}_{1}(t)$ is the residual in the singular basis.
+Averaging over samples, therefore, gives an error covariance $\bar{E}(t)\bar{E}(t)^{\top}$, and the continuous Oja update can be written as
+$\tau\frac{d\bar{P}}{dt}=\bar{P}\,\bar{E}\bar{E}^{\top}\,(I-\bar{P}^{\top}\bar{P})-\lambda\bar{P}.$ (16)
+Equations (5) and (16) thus define a coupled first-order dynamical system for $(\bar{W}_{1},\bar{W}_{2},\bar{P})$.
+This system has a straightforward interpretation.
+The Oja flow (16) steers the transmitted subspace toward the leading eigenvectors of the current error covariance, so the rank-$r$ channel allocates its constrained capacity to the modes that dominate the residual.
+As learning progresses, this mechanism tracks the dominant modes of the current residual, which in this solvable setting biases learning toward the leading singular directions of $\Sigma_{io}$ (Figure 4a).
+Figure 4: Low-dimensional feedback alignment (LDFA) as a synaptically local error pathway.
+(a) Linear-network analysis ($n=128$, $k=64$, $m=64$) with factorized feedback $B=QP$.
+Solid curves show simulations and dashed curves show the analytical prediction for the evolution of mode-wise singular values.
+With full-rank feedback ($r=m$), all task modes are recovered, whereas with rank-constrained feedback ($r<m$) learning prioritizes the dominant task modes selected by the Oja dynamics.
+(b) Multilayer perceptrons on CIFAR.
+Top: separating feedforward capacity from feedback dimensionality.
+A wide MLP reaches high accuracy even when the teaching signal is strongly compressed (LDFA with low rank $r$), whereas a much narrower MLP performs substantially worse despite having an unconstrained (full-rank) error pathway (BP).
+Bottom: varying task output dimensionality by subsampling classes from CIFAR-100 shows that the minimal rank required to match BP scales primarily with task dimension (number of classes), not with hidden-layer width (all layers use the same rank $r$).
+(c) Direct LDFA implements direct error projection by broadcasting a low-dimensional teaching signal to each layer either from the output layer (solid) or the penultimate layer (dashed), approaching BP-level accuracy as rank increases.
+(d) Convolutional networks trained with LDFA.
+A 4-block VGG-style CNN attains BP-level performance with low backward rank even when the widest blocks contain 512 channels.
+Inset: constraining all layers to a fixed fraction of their width ($r=n/2,n/4,n/8$) yields graceful degradation.
+5.2 Deep nonlinear networks
+We next test whether the local rule extends beyond the linear setting to deep nonlinear networks.
+Following the same procedure in section˜4, we train MLPs on CIFAR using the same constrained feedback architecture as above, where each layer receives an $r$-dimensional error signal through a fixed random mixing $Q_{l}$, while the projection $P_{l}$ is learned locally by the error-driven Oja update (15).
+Across architectures, we find that as with the normative approach in section˜3, rank required to match backpropagation tracks task dimensionality rather than network size.
+Concretely, when we subsample CIFAR-100 to classification problems with 50, 65 or 100 classes and constrain all feedback matrices to rank $r$, performance approaches full backpropagation once $r$ reaches the output dimension defined by the number of classes (Fig.
+4b).
+5.3 Low-dimensional error broadcasting
+A key limitation of the normative feedback-learning rule in section˜3 is its explicit dependence on the forward weights $W_{l}$, in the minimization of eq.˜8.
+This effectively assumes reciprocal connections between consecutive layers, and therefore cannot be applied to error pathways that bypass intermediate areas.
+In biological circuits, however, instructive signals are known to broadcast through indirect and long-range projections, while a mechanism that rely on local reciprocals has yet been found.
+A canonical example is the dopaminergic reward-prediction-error signal, which is delivered to many downstream targets via widespread projections [45, 62, 65].
+Direct feedback alignment (DFA) was introduced precisely to relax the requirement of reciprocal connectivity by broadcasting the output error directly to each hidden layer through fixed random feedback matrices [46].
+In its simplest form, each layer receives a teaching signal computed in parallel from the output-layer error, enabling updates without sequential backward transport through intermediate layers.
+While DFA can train deep networks in some regimes, its naive formulation typically exhibits a substantial performance gap relative to backpropagation on modern vision benchmarks, and convolutional architectures are a particularly challenging case [4, 14].
+Our local rule removes the architectural constraint that prevented the normative approach from operating in this regime.
+Concretely, for each layer $l$ we broadcast the error from a chosen source layer $l^{\prime}$ (often $L$) directly to layer $l$ using a low-rank matrix,
+$\boldsymbol{\delta}_{l}=\big(Q_{ll^{\prime}}P_{ll^{\prime}}\boldsymbol{\delta}_{l^{\prime}}\big)\odot f^{\prime}(a_{l}),\qquad l<l^{\prime}\leq L$ (17)
+where $l^{\prime}$ is an efferent layer to $l$ or the output error $\delta_{L}$.
+Same as above, $Q_{ll^{\prime}}$ is fixed and random and $P_{ll^{\prime}}\in\mathbb{R}^{r\times n_{l^{\prime}}}$ is learned by the local Oja update eq.˜15 (with $\boldsymbol{\delta}_{l^{\prime}}$ as its driving signal).
+This construction preserves the defining feature of DFA, namely parallel error broadcast through non-reciprocal pathways, while endowing the broadcast channel with an adaptive, low-rank projection that allocates its limited capacity to the dominant directions of the task error.
+Because the update of $P_{ll^{\prime}}$ is unsupervised and depends only on the locally available teaching signal, it applies unchanged when the feedback edges skip layers or originate from alternative sources (e.g., the penultimate layer), and it does not require access to the forward weights.
+Empirically, this direct error projection closes much of the gap between DFA and backpropagation in deep nonlinear networks.
+In particular, learning $P_{l}$ with $Q_{l}$ fixed suffices to train deep MLPs and CNNs with direct error broadcasts to backpropagation-level performance (Fig.
+4c), including regimes in which standard DFA is known to struggle.
+This demonstrates that the local subspace-learning mechanism is compatible with non-mirrored, indirect error pathways, strengthening the biological motivation of the framework while retaining the benefits of low-dimensional feedback.
+5.4 Convolutional networks
+A particularly stringent test is the convolutional setting.
+In deep CNNs, naive alignment schemes often exhibit a substantial gap relative to backpropagation unless augmented with additional mechanisms that improve alignment or stabilize training.
+Our low-dimensional feedback framework remains effective in this regime: we train CNNs reliably and reach backpropagation-level accuracy at comparable rank (Fig.
+3b), despite the combined challenges of deep credit assignment, weight sharing, and a restricted error channel.
+In all CNN experiments we learn the error projection $P_{l}$ using the local Oja update (15), which is the core mechanism that allocates the rank-$r$ channel to task-relevant error directions.
+Convolutional architectures, however, are a particularly stringent regime for alignment-based learning, and prior work on feedback alignment has typically found that allowing some adaptation of the feedback weights substantially improves robustness and accuracy in deep CNNs [34, 1].
+Following this established practice, we therefore also allow the mixing matrix $Q_{l}$ to adapt via a simple synaptically local Hebbian rule,
+$\Delta Q_{l}^{\mu}=\eta\,\boldsymbol{h}_{l}^{\mu}\big(P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\big)^{\top}-\lambda Q_{l},$ (18)
+which preserves the low-rank structure and does not introduce weight transport or any dependence on forward parameters.
+Importantly, (18) remains local: the presynaptic activity $\boldsymbol{h}_{l}^{\mu}$ and the postsynaptic teaching signal $P_{l}\boldsymbol{\delta}_{l+1}^{\mu}$ are available at the feedback synapse, and the update depends only on these signals and the synaptic weight.
+Empirically, this augmentation is consistent with the FA literature on challenging vision models and improves performance in convolutional architectures while retaining the central feature of LDFA: learning a low-dimensional error subspace through $P_{l}$.
+Figure 4d shows that training CNNs with low-dimensional error and local-only learning rules can match backpropagation performance, while maintaining more biological integrity.
 6.  Error dimensionality shapes neural receptive fields
-We have shown that neural networks can be efficiently trained using minimal error signals comparable to task dimensionality.
-Here, we investigate how error dimensionality affects neural representations, providing insights into receptive fields observed in the brain.
+Figure 5: Receptive fields are shaped by feedback dimensionality.
+(a) Simplified model of the early ventral visual stream (retina $\rightarrow$ VVS) adapted from [40], and two manipulations that decouple feedforward bottlenecks from the teaching pathway.
+Top: an anatomical bottleneck limits the retinal output.
+Middle: the feedforward architecture is left full-width, but the teaching signal to the retinal layer is restricted by a low-rank feedback map (LDFA).
+Bottom: the feedforward bottleneck is reinstated, while the retinal layer receives a high-dimensional teaching signal via direct error projection (dLDFA; Eq.
+(16)), bypassing backward compression.
+(b) Backpropagation reproduces the Lindsey et al.
+phenomenology: removing the retinal bottleneck yields oriented, edge-like receptive fields, whereas a narrow retinal bottleneck yields center-surround receptive fields.
+(c) With the feedforward pathway unconstrained but feedback to the retina rank-limited, LDFA reliably produces center-surround receptive fields; decreasing the feedback rank $r$ (examples $r=32,4,2$) increases their effective rotational symmetry.
+(d) Conversely, with a feedforward retinal bottleneck but high-dimensional error delivered directly to the retina using direct LDFA (dLDFA), receptive fields become oriented.
+Together, these manipulations show that the dimensionality and routing of the teaching signal can reproduce, and even override, receptive-field structure previously attributed to feedforward anatomical constraints.
+At the heart of our framework is the ability to decouple the error-feedback pathway from the feedforward pass, and to control the feedback channels independently.
+In Secs.
+3 and 4 we used this decoupling to study how restricting the backward pathway controls the efficiency of learning, and in Sec.
+5 we derived local update rules (including convolutional networks) that implement these constraints using synapse-local information.
+Here we ask a complementary question: beyond efficiency, does the structure of the error pathway shape the representations learned in early layers?
+Concretely, we study how receptive fields of shallow layers change when we constrain the dimensionality of the error signal reaching them.
+The relation between architecture and learned receptive fields has been studied before.
 Lindsey et al.
-found that in convolutional models of the visual system, narrow feedforward bottlenecks between the retina and the brain led to center-surround receptive fields in the retinal layer, similar to mammals [29].
-Wider bottlenecks resulted in orientation-selective receptive fields, as seen in salamanders.
-We hypothesize that these effects are due to constraints on the error signal reaching the retina rather than the feedforward bottlenecks themselves.
-To test this hypothesis, we trained a model similar to that of Lindsey et al.
-[29], but with full-width layers throughout.
-Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with LDFA (Figure 4c), thereby isolating the impact of error dimensionality on neural representations.
-We extracted the receptive fields of neurons in the retinal layer using visualization techniques [30] (Figure 4d).
-Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer.
-To validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dLDFA to train the retinal layer without restricting the backward pathway.
-In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix).
-This experiment demonstrates that error dimensionality influences neuronal tuning and neural representations.
-Specifically, lower-dimensional error signals promote higher symmetries in emergent receptive fields.
-Importantly, the architecture studied by Lindsey et al.
-is not a faithful model of the early visual system.
-In particular, there is no evidence that error signals propagate from the cortex or thalamus back to the retina.
-Instead, it is believed that the receptive fields in the retina emerge through unsupervised learning [31].
-Nevertheless, our results provide an explanation for the receptive fields observed by Lindsey et al.
-Furthermore, it may lead to an alternative explanation of learning in the visual pathway that uses simplified signaling (e.g., neuromodulators).
-Most importantly, our findings show a relation between tuning curves and error pathways.
+showed that in anatomically constrained convolutional models, narrowing the retinal output bottleneck yields center-surround receptive fields in the retinal layer, whereas widening it yields orientation-selective receptive fields [40].
+We hypothesize that this transition is driven primarily by the dimensionality of the teaching signal available to the retinal layer.
+With constrained feedback channels, the compressed error signal biases learning toward higher-symmetry solutions, even when the feedforward pathway is left unconstrained.
+LDFA provides a direct way to test this hypothesis because it lets us constrain the backward pathway while keeping the forward architecture fixed, and, via direct error projection (Eq.
+(17)), to manipulate error delivery independently of feedforward bottlenecks.
+We first remove the feedforward bottleneck by using full-width layers throughout, and restrict only the backward pathway by training a rank-$r$ feedback matrix to the retinal layer using LDFA (Fig.
+5a).
+We then visualize the resulting retinal receptive fields using standard feature-visualization methods [20] (Fig.
+5b,c).
+Despite the unconstrained feedforward architecture, restricting the feedback rank reliably produces center-surround receptive fields, and decreasing $r$ increases their effective rotational symmetry.
+We next perform the complementary manipulation.
+We reinstate a narrow feedforward bottleneck as in [40], but ensure that the retinal layer receives a high-dimensional error signal by using direct error projections (dLDFA; Eq.
+(17)), thereby bypassing any backward compression (Fig.
+5a bottom).
+In this setting, the retinal receptive fields become orientation selective (Fig.
+5d), matching the wide-bottleneck regime of [40].
+Together, these experiments show that the feedback pathway can reproduce, and even override, receptive-field phenomenology previously attributed to feedforward anatomical constraints and that error dimensionality acts as a powerful inductive bias on early-layer selectivity.
+The Lindsey et al.
+architecture is intentionally simplified and should not be read as a mechanistic model of retinal development; in particular, there is no evidence for supervised error signals propagating from cortex or thalamus back to the retina, and early retinal representations are widely thought to be shaped largely by local and unsupervised mechanisms [72].
+Our claim is therefore methodological and conceptual.
+Within a framework in which forward and backward passes are decoupled and implementable with local rules, the dimensionality and routing of learning signals can act as an inductive bias on receptive-field structure.
+This perspective applies most directly to early sensory stages that plausibly receive low-dimensional instructive signals (e.g., via neuromodulation or top-down pathways), and suggests that constraints on the error pathway may shape emergent selectivity at least as strongly as constraints on the feedforward circuit itself.
+This complements goal-driven modeling of sensory cortex by highlighting the backward pathway as an additional axis of inductive bias, alongside architecture and objective [70].
 7.  Discussion
-Our work demonstrates that neural networks can be trained effectively using minimal error signals constrained to the task’s intrinsic dimensionality rather than the higher dimensionality of the network’s representations.
-By adopting a factorized version of Feedback Alignment with low-rank matrices and training both left and right spaces of the feedback matrix, we showed that deep networks—linear, nonlinear, and convolutional—can match the performance of full backpropagation even under stringent error-dimensionality constraints.
-This finding highlights that the essential information required for learning is tied to the complexity of the task, measured by the number of output classes, and does not scale with the number of parameters or size of the model.
-Additionally, we revealed that constraining error dimensionality influences neural representations, providing a potential explanation for biological phenomena, such as center-surround receptive fields in the retina.
-This work aims to explore potential mechanisms for implementing gradient descent in the brain.
-Recognizing that high-dimensional feedback is not necessary for effective learning is a significant step toward developing more flexible and biologically realistic models of learning.
-This insight suggests that the brain might utilize low-dimensional error signals to drive learning processes, aligning with anatomical and physiological constraints.
+Our central result is that deep networks do not require transmitting the full layerwise gradient: it is sufficient to deliver a low-dimensional teaching signal to each layer, provided the error is projected into task-relevant directions.
+In a solvable linear theory, fixed low-rank feedback generically makes the stationarity conditions underdetermined and introduces spurious fixed points.
+Learning the feedback factors aligns the transmitted subspace with the evolving error and restores backpropagation-like convergence across nonlinear networks, including CNNs and vision transformers.
+This perspective cleanly separates a normative question (which error subspace should be communicated) from an implementational one (how that subspace can be learned).
+We first derive feedback learning as a low-rank factorization objective whose updates do not scale with batch size, and then show how essentially the same subspace can be learned with synapse-local plasticity driven by the arriving error, enabling direct error projection through non-mirrored pathways and revealing error dimensionality as an inductive bias that can shape early receptive fields.
+Computational efficiency.
+Because low-rank feedback constrains the error channel itself, the expensive multiplication by $W_{l}^{\top}$ in the backward pass is replaced by two thin multiplies, reducing per-example cost from $O(n_{l}n_{l+1})$ to $O(r(n_{l}+n_{l+1}))$ and making backward FLOPs scale linearly with the effective error dimensionality $r$.
+End-to-end, this creates a trade-off in which decreasing the rank $r$ lowers computational cost, but excessive compression degrades learning efficiency, requiring more samples.
+Consequently, total compute exhibits a U-shaped dependence on $r$, with an intermediate rank minimizing FLOPs-to-accuracy in our transformer experiments.
+This positions low-dimensional feedback as a structured alternative to gradient sparsification and compression methods such as meProp [66] and Deep Gradient Compression [39].
+Rather than sparsifying a full gradient after it is formed, we learn the projection that determines which error directions are represented upstream, so the savings come from never materializing or transporting the discarded components.
+Aligning to the error space.
+At first glance, our local rule resembles adaptive variants of Feedback Alignment (FA), where the feedback weights are trained to better match the feedforward weights and thereby approximate backpropagation [38, 1].
+The factorization $B_{l}=Q_{l}P_{l}$ reveals a different target: $Q_{l}$ only mixes an $r$-dimensional teaching channel into layer $l$, whereas $P_{l}$ determines which output-error directions can be represented and transmitted upstream (the subspace spanned by $P_{l}^{\top}$).
+Learning $P_{l}$ therefore aligns the transmitted error subspace with the dominant modes of the current residual, improving the information content of the teaching signal itself.
+This perspective explains why fixed low-rank feedback can induce spurious fixed points, and why subspace learning resolves them: the bottleneck fails when it discards task-relevant error directions, not merely when it misaligns with $W_{l}^{\top}$.
+Low-rank adaptation (LoRA) in large language models.
+Recent parameter-efficient fine-tuning methods for large language models, most notably LoRA, achieve strong adaptation by freezing pretrained weights and learning low-rank updates $\Delta W$ within transformer layers [26], with practical extensions that push memory efficiency further (e.g., quantization with QLoRA) and adaptive rank allocation) [17, 41].
+Our approach is complementary in where the low-rank constraint is imposed.
+Here, instead of restricting the forward update parameterization while retaining full backpropagated errors, we restrict and learn the error subspace communicated to each layer.
+This provides a middle ground in which the number of trainable forward parameters can grow (increasing representational capacity) while the cost of credit assignment remains controlled by the feedback rank.
 Bridging biological constraints and gradient-based learning.
-Our work aims to reconcile gradient-based learning with the anatomical constraints of cortical circuits.
-In the cortex, feedback connections are generally sparser, less structured, and more selective than their feedforward counterparts [32], suggesting that learning may rely on a restricted top-down error signal.
-Several alternative frameworks have been proposed to explain learning in hierarchical biological circuits.
-Predictive coding models suggest that the cortex continuously generates predictions about sensory input and computes discrepancies, or prediction errors, between expected and actual signals [33, 34] .
-Equilibrium propagation [35] offers another biologically inspired alternative, demonstrating how energy-based models can approximate backpropagation through local weight updates.
-While these approaches provide valuable insights into neural computation, they deviate from the conventional framework of gradient-based optimization.
-In contrast, our approach remains firmly within the gradient descent paradigm, allowing direct loss minimization.
-This key property enables its seamless application across diverse architectures, from fully connected nonlinear networks to convolutional models and transformer-based attention mechanisms.
-Between reward-based learning and full gradient descent.
-By showing that deep networks can be trained with low-dimensional error signals, our framework bridges the gap between gradient-based learning and reward-based learning.
-Studies have suggested that dopaminergic neurons in the ventral tegmental area (VTA) encode reward signals that capture multiple task attributes rather than a single scalar value [36, 37, 38].
-Our findings provide a theoretical foundation for moving beyond scalar reward-based learning toward a framework where low-dimensional reward signals carry richer information about the underlying loss (or gain).
-This perspective positions our model as an intermediary between reinforcement learning’s scalar feedback and the full backpropagation of error gradients.
-More broadly, our results suggest that learning could emerge through the strategic compression of error signals into task-relevant representations, offering a biologically plausible alternative to high-dimensional gradient propagation.
-Aligning to the error space.
-On its surface, our novel learning rule is similar to previous Feedback Alignment (FA) schemes.
-However, it offers a conceptual novelty.
-In traditional FA, learning the feedback weights aims to align them with the feedforward weights [5, 10], mirroring full backpropagation.
-In contrast, by factorizing the feedback matrix $B=QP$, we also align the row space of the feedback matrix with the source of the error, thereby improving the quality of the error signal itself.
-Computational complexity.
-Despite the additional requirement of training the feedback matrices, the computational cost of using low-rank feedback does not increase significantly.
-The updates for the low-rank matrices are linear in the number of neurons or filters per layer, while the forward pass and the primary weight updates remain quadratic [39].
-This efficiency makes our approach practical for large-scale networks and suggests that similar mechanisms could be feasible in biological neural networks.
-Gradient descent in high dimensions.
-Our findings invite a rethinking of gradient descent dynamics in large, overparameterized networks.
-Typically, the weight dynamics during training are high-dimensional [40].
-However, when the error signal is low-dimensional, the weight updates in each layer are confined to a much lower-dimensional subspace.
-This constraint could have implications for understanding the generalization capabilities of neural networks, as it suggests that effective learning does not require exploring the full parameter space.
-Indeed, previous studies found that the Hessian of the loss function shows low-rank structure [41].
-Exploring this connection could open new avenues for understanding the dynamics of gradient descent in high-dimensional loss spaces [42].
-Exploring the potential of low-dimensional feedback in more complex architectures.
-This approach not only improves learning efficiency but may also act as a form of regularization, as can be seen from the improved performance of low-rank feedback matrices on visual transformers—the possibility of using restricted error signals as a form of regularization warrants further investigation.
-In summary, our findings not only challenge the conventional reliance on high-dimensional error propagation but also pave the way for the development of biologically inspired, computationally efficient learning algorithms that hold promise for advancing both artificial intelligence and our understanding of brain function." This would leave the reader with a clear impression of both the scientific and translational impact of your work
+Our work addresses a basic anatomical mismatch between backpropagation and cortex: feedback pathways are not organized as a synapse-wise transpose of the forward weights and need not deliver a neuron-by-neuron teaching signal [43].
+Predictive coding and equilibrium propagation propose mechanistic routes by which local plasticity, together with inference dynamics, can approximate gradient-based learning in hierarchical circuits [53, 69, 61].
+Our contribution is complementary: we remain in the standard objective-minimization setting, but treat the teaching pathway as a constrained resource and learn a low-dimensional projection of error signals that preserves task-relevant directions.
+This yields effective credit assignment under severe bandwidth limits and, via direct error projection, remains compatible with indirect and non-reciprocal delivery of instructive signals.
+Low-dimensional teaching signals and reward pathways.
+The same perspective offers a middle ground between scalar reinforcement signals and full gradient transport.
+Midbrain dopamine has long been linked to reward-prediction-error teaching signals, yet recordings increasingly suggest heterogeneous coding in which dopaminergic populations and terminals reflect multiple sensory, motor, and cognitive variables beyond a single scalar reward [13, 49, 19].
+In our framework, such signals can be interpreted as a low-dimensional teaching vector: a compressed, broadcast error representation whose efficacy depends on learning the appropriate projection into each layer’s plastic synapses.
+This suggests a concrete computational role for low-dimensional neuromodulatory feedback, without requiring the brain to propagate a full high-dimensional gradient through deep circuits.
+Feedback structure as an inductive bias.
+Inductive biases in deep learning are usually discussed as properties of the forward architecture, where weight sharing, locality, and symmetry constraints bias learning toward particular solution classes and improve generalization [71, 8, 64, 23].
+Our results suggest that the teaching channel can play a similar role by constraining and shaping the feedback pathway.
+The feedback path controls which error directions are expressible upstream, thereby biasing optimization toward representations that are easiest to learn under the available error bandwidth.
+This viewpoint is consistent with recent analyses showing that feedback-alignment-style learning exhibits implicit regularization and can select among degenerate low-loss solutions [54].
+In our setting, the receptive-field experiments provide a concrete illustration that modifying only the feedback pathway can systematically alter learned symmetries.
+A promising direction is therefore to treat feedback design as a new knob for algorithmic inductive bias: for example, imposing topographic, convolutional, modular, or sign-constrained structure on low-rank feedback may preferentially promote invariances, robustness, or task-specific feature reuse, while maintaining computational efficiency through a controlled error bottleneck.
+In summary, we argue that the essential constraint for scalable credit assignment is not access to a full, layerwise gradient but access to a learned low-dimensional teaching channel that preserves task-relevant error directions.
+This reframes feedback as a resource that can be optimized, offering a principled route to more efficient training when backward computation or communication is limiting, and a concrete hypothesis for how anatomically restricted pathways could still support deep learning.
+More broadly, our results invite future work that treats the structure of feedback, not only the forward architecture, as a design axis that shapes both efficiency and the inductive biases of learned representations.
 Acknowledgments
-This work was supported by the Gatsby Charitable Foundation and the Azrieli Foundation.
+This work was supported by the Gatsby Charitable Foundation.
 Appendix
 Appendix A Full linear theory
-in this section, we provide the analysis of the linear networks for both Feedback Alignment and LDFA
+In this section, we provide the analysis of the linear networks for both Feedback Alignment and LDFA.
 Detailed analysis of Feedback Alignment learning dynamics
@@
 $\tau\frac{d\bar{W}_{2}}{dt}$ $=\left(S-\bar{W}_{2}\bar{W}_{1}\right)\bar{W}_{1}^{\top}.$
-Full derivation of low-dimensional feedback alignment (LDFA) learning dynamics
+Derivation of Normative Low-Rank Feedback Learning
+Here we derive the learning dynamics for the normative variant of LDFA, where the feedback matrix $B=QP$ is trained to explicitly approximate the transpose of the forward weights $W_{2}^{\top}$.
+The feedback factors $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$ are updated to minimize the reconstruction loss:
+$\mathcal{L}_{B}=\frac{1}{2}\|QP-W_{2}^{\top}\|_{F}^{2}.$ (A27)
+The gradients with respect to the feedback factors are:
+$\frac{\partial\mathcal{L}_{B}}{\partial P}$ $=Q^{\top}(QP-W_{2}^{\top}),$ (A28)
+$\frac{\partial\mathcal{L}_{B}}{\partial Q}$ $=(QP-W_{2}^{\top})P^{\top}.$
+Assuming gradient descent with a time constant $\tau_{B}$, the continuous-time dynamics are:
+$\tau_{B}\frac{dP}{dt}$ $=Q^{\top}(W_{2}^{\top}-QP),$ (A29)
+$\tau_{B}\frac{dQ}{dt}$ $=(W_{2}^{\top}-QP)P^{\top}.$
+We now rotate these dynamics into the singular basis of $\Sigma_{io}$ using the same transformations as above: $W_{2}=U\bar{W}_{2}$ (implying $W_{2}^{\top}=\bar{W}_{2}^{\top}U^{\top}$) and $P=\bar{P}U^{\top}$.
+Note that $Q$ operates in the hidden space and is not rotated by $U$.
+Substituting these into the dynamics for $P$:
+$\tau_{B}\frac{d(\bar{P}U^{\top})}{dt}$ $=Q^{\top}\left(\bar{W}_{2}^{\top}U^{\top}-Q\bar{P}U^{\top}\right)$ (A30)
+$=Q^{\top}\left(\bar{W}_{2}^{\top}-Q\bar{P}\right)U^{\top}.$ (A31)
+Multiplying by $U$ on the right yields the rotated dynamics for $\bar{P}$:
+$\tau_{B}\frac{d\bar{P}}{dt}=Q^{\top}\left(\bar{W}_{2}^{\top}-Q\bar{P}\right).$ (A32)
+Similarly for $Q$:
+$\tau_{B}\frac{dQ}{dt}$ $=\left(\bar{W}_{2}^{\top}U^{\top}-Q\bar{P}U^{\top}\right)(\bar{P}U^{\top})^{\top}$ (A33)
+$=\left(\bar{W}_{2}^{\top}-Q\bar{P}\right)U^{\top}U\bar{P}^{\top}$ (A34)
+$=\left(\bar{W}_{2}^{\top}-Q\bar{P}\right)\bar{P}^{\top}.$ (A35)
+Together with the forward weight updates derived in Eq. (A19) and (A25), this forms the closed dynamical system described in Eq. (10) of the main text.
+Full derivation of local low-dimensional feedback alignment (LDFA) learning dynamics
 We consider an alternative algorithm where the matrix $B$ is replaced by $B=QP$, with $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$.
@@
 $\tau\frac{dQ}{dt}=\bar{W}_{1}V_{1}^{\top}\left(USV^{\top}-U\bar{W}_{2}\bar{W}_{1}V^{\top}\right)^{\top}(PU^{\top})^{\top}$ (A47)
-$=\bar{W}_{1}V_{1}^{\top}V\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{T}U^{\top}U\bar{P}^{\top}$ (A39)
+$=\bar{W}_{1}V_{1}^{\top}V\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}U^{\top}U\bar{P}^{\top}$ (A48)
 $=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$ (A49)
@@
 $\tau\frac{d\bar{P}}{dt}$ $=\bar{P}S^{2}\left(I-\bar{P}^{\top}\bar{P}\right).$
-$\tau\frac{d\bar{Q}}{dt}$ $=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$
+$\tau\frac{dQ}{dt}$ $=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$
 Appendix B Learning rule implementation
@@
 For the LDFA layers, we initialized all weights using Kaiming uniform initialization.
-We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule.
-In the output layer, we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$.
-In the hidden layers, $\boldsymbol{P}$ is learned to project onto the principal directions of the error signal from the subsequent layer, represented as $\boldsymbol{\delta}_{l+1}$.
+We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule as well as calculating the gradients of $P$ and $Q$.
+We implemented two variants of the algorithm.
+In the **Normative** version (Algorithm 1), $P$ and $Q$ are updated to explicitly minimize the reconstruction error $\|QP-W^{\top}\|$, ensuring the feedback pathway approximates the transpose of the forward weights.
+In the **Local** version (Algorithm 2), $P$ and $Q$ are updated using biologically plausible, local plasticity rules (Oja’s rule) that align the feedback subspace with the principal components of the error signal.
+Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$,
+     $P_{l}\in\mathbb{R}^{r\times m}$, weight $W_{l}\in\mathbb{R}^{m\times n}$
+Output: Gradient w.r.t. input grad_input, updates $\Delta Q_{l}$, $\Delta P_{l}$, $\Delta W_{l}$
+Compute ’gradient’ w.r.t. input:
+$\text{grad\_input}\leftarrow\left(Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\top}\right)^{\top}\in\mathbb{R}^{b\times n}$;
+Compute update for $W_{l}$:
+$\Delta W_{l}\leftarrow\boldsymbol{\delta}_{l+1}^{\top}\boldsymbol{h}_{l}$;
+Compute updates for $Q_{l}$ and $P_{l}$ (Reconstruction):
+$E\leftarrow W_{l}^{\top}-Q_{l}P_{l}$;
+$\Delta Q_{l}\leftarrow EP_{l}^{\top}$;
+$\Delta P_{l}\leftarrow Q_{l}^{\top}E$;
+Algorithm 1 Backward Pass for Normative LDFA (Approximating $W^{\top}$)
+In the local version (Algorithm 2), we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$ (in the output layer) or the error signal $\boldsymbol{\delta}_{l+1}$ (in hidden layers).
 Specifically, for a layer $l$ with input dimension $n$, output dimension $m$, and rank constraint $r$, we proceed as follows:
-Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$, $P_{l}\in\mathbb{R}^{r\times m}$, weight matrix $W_{l}\in\mathbb{R}^{m\times n}$
+Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$,
+     $P_{l}\in\mathbb{R}^{r\times m}$, weight $W_{l}\in\mathbb{R}^{m\times n}$
 Output: Gradient w.r.t. input grad_input, updates $\Delta Q_{l}$, $\Delta P_{l}$, $\Delta W_{l}$
-Compute covariance matrix:
+Compute ’gradient’ w.r.t. input:
+$\text{grad\_input}\leftarrow\left(Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\top}\right)^{\top}\in\mathbb{R}^{b\times n}$;
+Compute update for $W_{l}$:
+$\Delta W_{l}\leftarrow\boldsymbol{\delta}_{l+1}^{\top}\boldsymbol{h}_{l}$;
+Compute covariance matrix for P update:
 if use_targets then
@@
 $\gamma\leftarrow\displaystyle\max\left(\operatorname{diag}\left(C\right)\right)$;
-Compute ’gradient’ w.r.t. input:
-[…]
-Compute updates for $Q_{l}$ and $P_{l}$:
+Compute updates for $Q_{l}$ and $P_{l}$ (Local):
 $\Delta Q_{l}\leftarrow\boldsymbol{h}_{l}^{\top}\left(P_{l}\boldsymbol{\delta}_{l+1}^{\top}\right)^{\top}$;
 $\Delta P_{l}\leftarrow P_{l}\left(\dfrac{C}{\gamma}\right)\left(I-P_{l}^{\top}P_{l}\right)$;
-Compute update for $W_{l}$:
-$\Delta W_{l}\leftarrow\boldsymbol{\delta}_{l+1}^{\top}\boldsymbol{h}_{l}$;
-Algorithm 1 Modified Backward Pass for LDFA Layer
-The derivation of the dLDFA (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
+Algorithm 2 Backward Pass for Local LDFA (Oja’s Rule)
+The derivation of the dLDFA (directed restricted adaptive feedback) follows the same procedure as Algorithm 2, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
 Appendix C Numerical experiments
+C.1 Common Architectures and Training Details
+Across our experiments, we utilized three primary architectures: a fully connected Multi-Layer Perceptron (MLP), a VGG-style Convolutional Neural Network (CNN), and a Vision Transformer (ViT).
+Unless otherwise stated in the specific experiment subsections, the architectures follow the specifications below.
+Multi-Layer Perceptron (MLP)
+For the dense network experiments on CIFAR-10 and CIFAR-100, we used a network with 4 hidden layers.
+  • Structure: Input $\to$ [Linear(512) $\to$ ReLU] $\times$ 3 $\to$ Linear(Output).
+  • Hidden Width: 512 neurons (unless specified otherwise for scaling experiments).
+  • Initialization: Kaiming Uniform.
+Convolutional Neural Network (VGG-Like)
+We trained a VGG-style network with four convolutional blocks and batch normalization. The detailed architecture is provided in Table C1.
+Table C1: VGG-style CNN Architecture used for CIFAR-10 experiments.
+Block Layer Type Details
+Block 1 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 64 filters, padding=1
+MaxPool2d $2\times 2$ kernel, stride=2
+Output Size $16\times 16\times 64$
+Block 2 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 128 filters, padding=1
+MaxPool2d $2\times 2$ kernel, stride=2
+Output Size $8\times 8\times 128$
+Block 3 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 256 filters, padding=1
+MaxPool2d $2\times 2$ kernel, stride=2
+Output Size $4\times 4\times 256$
+Block 4 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 512 filters, padding=1
+AdaptiveAvgPool2d Output size $(1,1)$
+Flatten Vector size 512
+Classifier Linear + ReLU $512\to 256$
+Dropout $p=0.4$
+Linear $256\to\text{Num Classes}$
+Vision Transformer (ViT)
+We utilized the Vision Transformer implementation from the timm library. The default configuration used across experiments (unless explicitly modified for scaling tests) is detailed in Table C2.
+Table C2: Vision Transformer (ViT) hyperparameters.
+Parameter Value
+Patch Size 4
+Input Channels 3
+Embedding Dimension 384 (variable in scaling experiments)
+Depth 8 layers
+Number of Heads 8
+MLP Ratio 2.0
+QKV Bias True
+Drop Rate 0.1
+Attention Drop Rate 0.1
+Drop Path Rate 0.1
+C.2 Experiments with Normative Feedback
+In this section, we detail the experiments corresponding to Section 4 of the main text, where the feedback matrices $B_{l}=Q_{l}P_{l}$ are trained to minimize the reconstruction error $\|Q_{l}P_{l}-W_{l}^{\top}\|_{F}^{2}$ (Algorithm 1).
+Dual Optimization Strategy
+In all normative experiments, we employed two separate optimizers: one for the forward weights ($W_{l}$) and one for the backward feedback factors ($Q_{l},P_{l}$).
+Both used the AdamW algorithm but were configured with different learning rates and schedules to ensure the feedback pathway effectively tracked the evolving forward weights.
+Convolutional Networks (Layer-wise Constraints)
+To evaluate the impact of error compression on deep convolutional networks (Fig. 3a), we trained the VGG-like architecture defined in Table C1 on the CIFAR-10 dataset.
+  • Rank Constraints: We replaced the standard backward pass in every convolutional layer with a low-rank feedback channel.
+The rank $r$ was set to a fixed fraction of the layer’s channel width $n$ (i.e., $r\in\{n/2,n/4,n/8\}$).
+  • Feedback Kernel: For convolutional layers, the projection $P_{l}$ operates as a $1\times 1$ convolution, reducing the error signal from $m$ channels to $r$ channels per spatial location.
+Vision Transformers (ViT)
+We trained the Vision Transformer using the dual-optimizer setup described below. The training duration was set to 250 epochs for CIFAR-10 and 300 epochs for CIFAR-100.
+  • Scope: Low-rank feedback was applied to all learnable linear transformations, including Attention (Q, K, V) and MLP layers.
+  • Forward Optimization: The forward weights were optimized using AdamW with a base learning rate of $3\times 10^{-4}$ and weight decay of $0.1$.
+  • Backward Optimization: The feedback factors were optimized using AdamW with a base learning rate of $0.015$ and no weight decay.
+  • Scheduling: Both optimizers utilized a warmup phase of 10 epochs with a start factor of 0.1 (via linear scheduling), followed by a cosine annealing schedule decaying to a minimum learning rate of $1\times 10^{-9}$.
+  • CIFAR-100 Adaptation: For CIFAR-100 experiments, the backward optimizer’s base learning rate was adjusted to $0.005$.
+C.3 Experiments with Local Feedback
+In this section, we verify the efficacy of the local learning rule (Algorithm 2), where the feedback projection $\boldsymbol{P}_{l}$ is learned via an error-driven Oja’s rule to capture the principal subspace of the teaching signal.
+Multi-Layer Perceptron (MLP) Experiments
+Experimental Setup
+Unless otherwise noted, we used the standard MLP architecture described in the Common Architectures section (4 hidden layers, 512 neurons).
+All models were trained for 160 epochs using the Adam optimizer with AMSGrad, a batch size of 32, a learning rate of $6\times 10^{-4}$, and a weight decay of $4\times 10^{-4}$.
+The learning rate followed an exponential decay with a factor of 0.975.
 Layer-Wise Constraints
-In Figure [3.a], we present the results of training a network with four hidden layers, each containing 512 neurons, and an output layer with 10 neurons on the CIFAR-10 dataset.
-The network was trained using LDFA, as described earlier, without rank constraints, except for one layer at a time.
-For comparison, we also trained the network using standard backpropagation as a baseline.
-All networks were trained with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$.
-The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975.
+To analyze the sensitivity of different depths to error compression, we applied the LDFA rank constraint to only one layer at a time, leaving all other layers unconstrained.
+We compared these distinct training runs against a baseline trained with standard backpropagation.
 Each experiment was repeated 10 times.
 Constraining All Layers
-For the results shown in Figure [3.b], we trained the same network with four hidden layers, each containing 512 neurons.
-This time, we applied rank constraints to all layers simultaneously, with rank values $r=64,32,16,10$.
-To further demonstrate that the network still utilizes high-rank representations, we also trained a variant of the network with 64 neurons in each hidden layer, without applying any rank constraints.
-All training was conducted with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$.
-The Adam optimizer with AMSGrad was used, with training carried out for 160 epochs and an exponential learning rate decay factor of 0.975.
-Each experiment was repeated 10 times.
-CIFAR-100 Sub-sampling
-For the results shown in Figure [3.c], we trained the same model on the CIFAR-100 dataset, sampling different numbers of classes $d$, with $d=50,75,100$.
-For each sub-sample, we trained the model while applying rank constraints to all layers, using various rank values.
-This was done to demonstrate that the dimensionality of the error signal depends on the task dimensionality $d$.
-All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$.
-The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975.
-Each training run was repeated 5 times.
-Direct LDFA
-For the results shown in Figure [3.d], we trained the same model using dLDFA.
-In one experiment, we propagated the error signal from the last layer to all preceding layers.
-In a separate experiment, we propagated the error signal directly from the penultimate layer to all earlier layers, applying rank constraints to these layers while keeping the last layer at full rank.
-All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$.
-The Adam optimizer with AMSGrad was used, training for 160 epochs with an exponential learning rate decay factor of 0.975.
-Each training run was repeated 5 times
-Convolutional Neural Networks
-To extend our LDFA algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal.
-This effectively constrains the dimensionality of the error signal across the spatial dimensions.
-Similar to the implementation for fully connected layers in Algorithm 1, we modify the backward pass in the same way for convolutional layers.
-The key difference in the convolutional context is how the projection matrix $\boldsymbol{P}$ operates on the error signals.
-In convolutional layers, $\boldsymbol{P}$ functions as a $1\times 1$ convolutional filter, projecting the error signal at each spatial location from $m$ channels down to $r$ channels.
-This reduces the error signal’s dimensionality to $r$ per pixel, adhering to the rank constraint.
-As in the fully connected case, the matrix $\boldsymbol{P}$ is learned using Oja’s rule.
-However, the covariance matrix $\boldsymbol{C}$ is computed over both the batch and spatial dimensions—that is, across all pixels in all images within the batch.
-This approach captures the covariance structure of pixel representations more effectively, enabling $\boldsymbol{P}$ to project the error signals appropriately in the convolutional setting.
-VGG-like Architecture
-We trained a VGG-like convolutional neural network comprising four convolutional blocks, as detailed in Table C1.
-For the experiments presented in Figure 4(a), a rank constraint was applied exclusively to the final convolutional block, which consists of 512 channels.
-In contrast, for the results shown in Figure 4(b), rank constraints were imposed on all layers, reducing the rank to 1/2, 1/4, and 1/8 of the original channel dimensions.
-Each network was trained using the Adam optimizer with a learning rate of $5\times 10^{-4}$, a weight decay of $5\times 10^{-5}$, and an exponential learning rate decay factor of 0.98.
-Training was conducted for 250 epochs, with each experiment repeated five times to ensure robustness.
-Vision Transformer (ViT)
-We implemented a Vision Transformer (ViT) model that consists of seven self-attention blocks, each with an embedding dimension of 384 and four attention heads.
-The input images were partitioned into nonoverlapping patches, which were linearly projected into a 384-dimensional feature space.
-These patch embeddings were augmented with learned positional encodings and processed through a stack of seven transformer encoder layers.
-Each encoder block employed multi-head self-attention with four heads, followed by pre-norm layer normalization and dropout for regularization.
-The final representation was obtained from a learned class token, which was subsequently processed by a fully connected layer for classification.
-Training was carried out using a batch size of 32, a learning rate of $3\times 10^{-4}$, and a weight loss of $1\times 10^{-4}$.
-We used the Adam optimizer with AMSGrad, training for 150 epochs with an exponential learning rate decay factor of 0.98.
-Each experiment was repeated five times to ensure robustness.
+We evaluated the performance when the entire error pathway is compressed by applying rank constraints simultaneously to all layers with ranks $r\in\{64,32,16,10\}$.
+To verify that the network relies on high-dimensional feedforward representations even when feedback is low-rank, we trained a control model where the feedforward width was physically reduced to 64 neurons (matching the feedback bottleneck) without rank constraints.
+These experiments were repeated 10 times.
+Task Dimensionality (CIFAR-100)
+To test the hypothesis that feedback rank scales with task dimensionality $d$, we trained the model on CIFAR-100 subsets with $d\in\{50,75,100\}$ classes.
+For each subset, we applied rank constraints to all layers.
+Each run was repeated 5 times.
+Direct LDFA (dLDFA)
+We evaluated the local rule’s ability to support broadcast error signals in two configurations:
+  • Output Broadcast: The error signal $\boldsymbol{\delta}_{L}$ was propagated directly from the output layer to all preceding hidden layers.
+  • Penultimate Broadcast: The error signal was propagated from the penultimate layer to all earlier layers (applying rank constraints), while the final layer was updated normally.
+These runs were repeated 5 times.
+Convolutional Neural Networks (CNN)
+Methodology
+To extend LDFA to convolutional layers, we treat the rank constraint as a bottleneck on the number of error channels.
+The projection matrix $\boldsymbol{P}$ functions as a $1\times 1$ convolutional filter, projecting the error signal at each spatial location from $m$ channels down to $r$ channels.
+Crucially, the covariance matrix $C$ for Oja’s rule (Algorithm 2) is computed by aggregating over both the batch and spatial dimensions (i.e., across all pixels in all images).
+This ensures $\boldsymbol{P}$ captures the covariance structure of the pixel-wise error representations.
+VGG-like Experiments
+We trained the VGG-style architecture defined in Table C1.
+Training was conducted for 250 epochs using Adam with a learning rate of $5\times 10^{-4}$, weight decay of $5\times 10^{-5}$, and an exponential decay factor of 0.98.
+All experiments were repeated 5 times.
 Neural Receptive Fields
@@
 Additionally, we conducted experiments with dynamic Restricted Adaptive Feedback (dLDFA), where feedback to the retina originated from higher visual layers while feedforward connections within the retina were constrained to four channels.
-Table C1: CNN Architecture
-Layer(s) Output Size Details
-Input $32\times 32\times 3$
-Convolutional Block 1
-Conv2D + ReLU $32\times 32\times 64$ $3\times 3$ conv, 64 filters, padding=1
-BatchNorm2D $32\times 32\times 64$
-Conv2D + ReLU $32\times 32\times 64$ $3\times 3$ conv, 64 filters, padding=1
-BatchNorm2D $32\times 32\times 64$
-MaxPool2D $16\times 16\times 64$ $2\times 2$ max pool, stride=2
-Convolutional Block 2
-Conv2D + ReLU $16\times 16\times 128$ $3\times 3$ conv, 128 filters, padding=1
-BatchNorm2D $16\times 16\times 128$
-Conv2D + ReLU $16\times 16\times 128$ $3\times 3$ conv, 128 filters, padding=1
-BatchNorm2D $16\times 16\times 128$
-MaxPool2D $8\times 8\times 128$ $2\times 2$ max pool, stride=2
-Convolutional Block 3
-Conv2D + ReLU $8\times 8\times 256$ $3\times 3$ conv, 256 filters, padding=1
-BatchNorm2D $8\times 8\times 256$
-Conv2D + ReLU $8\times 8\times 256$ $3\times 3$ conv, 256 filters, padding=1
-BatchNorm2D $8\times 8\times 256$
-MaxPool2D $4\times 4\times 256$ $2\times 2$ max pool, stride=2
-Convolutional Block 4
-Conv2D + ReLU $4\times 4\times 512$ $3\times 3$ conv, 512 filters, padding=1
-BatchNorm2D $4\times 4\times 512$
-Conv2D + ReLU $4\times 4\times 512$ $3\times 3$ conv, 512 filters, padding=1
-BatchNorm2D $4\times 4\times 512$
-AdaptiveAvgPool2D $1\times 1\times 512$ Output size $(1,1)$
-Flatten $512$ Flatten to vector
-Classifier
-Fully Connected + ReLU $256$ Linear layer, $512\rightarrow 256$
-Dropout $256$ Dropout probability $p=0.4$
-Fully Connected $C$ Linear layer, $256\rightarrow C$
 Table C3: Retina model Architecture